### v1

# AutoRev: Automatic Peer Review System for Academic Research Papers

Abstract
Generating a review for an academic research paper is a complex task that requires a deep understanding of the document’s content and the interdependencies between its sections. It demands not only insight into technical details but also an appreciation of the paper’s overall coherence and structure. Recent methods have predominantly focused on fine-tuning large language models (LLMs) to address this challenge. However, they often overlook the computational and performance limitations imposed by long input token lengths. To address this, we introduce AutoRev, an Automatic Peer Review System for Academic Research Papers. Our novel framework represents an academic document as a graph, enabling the extraction of the most critical passages that contribute significantly to the review. This graph-based approach demonstrates effectiveness for review generation and is potentially adaptable to various downstream tasks, such as question answering, summarization, and document representation. When applied to review generation, our method outperforms SOTA baselines by an average of 58.72% across all evaluation metrics. We hope that our work will stimulate further research in applying graph-based extraction techniques to other downstream tasks in NLP. We plan to make our code public upon acceptance.
AutoRev: Automatic Peer Review System for Academic Research Papers
Maitreya Prafulla Chitale1thanks: Equal contribution.  Ketaki Mangesh Shetye [^1] footnotemark: 1 Harshit Gupta1Manav Chaudhary1Vasudeva Varma1 1IIIT Hyderabad {maitreya.chitale, ketaki.shetye, harshit.g, manav.chaudhary}@research.iiit.ac.in vv@iiit.ac.in
1 Introduction
Figure 1: Example graph constructed out of a paper consisting of 2 Headings, 2 Sub-Headings, 4 Passages, and 6 Sentences.
Recent research has increasingly focused on developing automatic systems for peer review, including tools that assist reviewers, assess peer-review processes, and directly generate reviews for academic papers Yu et al. (2024); Aitymbetov and Zorbas (2025); Church et al. (2025). Concurrently, there has been growing interest in examining the strengths and limitations of such systems, particularly the use of large language models (LLMs) for automated review generation Ye et al. (2024). Motivated by both of these trends, we propose AutoRev, a system designed to generate peer reviews from a guidance perspective, offering constructive feedback to support researchers and reviewers.
With the rise of Graph Neural Networks, the use of graphs as extractors or as guides for facilitating downstream tasks has recently gained traction in the NLP community Chitale et al. (2025); Gao et al. (2025); Agarwal et al. (2025). Inspired by this, we propose to use graphs in the domain of academic paper analysis, specifically, generating reviews that closely replicate those produced by existing peer review systems in conferences. AutoRev can generate high-quality reviews without relying on any prior human evaluations of the paper. Furthermore, by leveraging the graph’s capability as an extractor of passages, we significantly reduce the input token length to the LLM. Reducing the token length of the input has been shown to enhance LLM performance Hsieh et al. (2024).
To summarize our main contributions:
  • We propose a system for generating concise, structured reviews (Summary, Strengths, Weaknesses, and Questions) to assist authors and reviewers.
  • To our knowledge, this is the first approach using graphs in this domain.
  • Our approach improves over baselines by 58.72% on average across all metrics, without relying on human-written reviews and staying efficient.
It is important to highlight that this paper does not seek to determine the final reviews of a submission. Rather, we present our framework as a tool to provide timely feedback to researchers and serve as a valuable guide for reviewers.
2 Related Works
A significant amount of research has been conducted on aiding peer reviews for academic papers. Some studies explore the involvement of AI in the selection process Aitymbetov and Zorbas (2025), while some studies work on generating meta-reviews Kumar et al. (2023). Sufficient discussion and avenues have been explored where AI/NLP techniques can contribute towards aiding automatic peer reviews Kuznetsov et al. (2024). Though, to the best of our knowledge, significant research has not been done in the field of generating a whole review for an academic paper.
A recent study by Faizullah et al. (2024) explores a DPR-based approach for generating suggestive limitations. It highlights the necessity of using DPR to address length constraints and extract only the most relevant passages from a paper. This study served as an initial motivation for incorporating DPR into our approach.
Another notable work by Yu et al. (2024) introduces an automated paper reviewing framework. The work also includes a standardized dataset of GPT-4 OpenAI et al. (2024) consolidated reviews collected from various academic venues. To enhance review generation, the authors propose a self-correction strategy during LLM fine-tuning to ensure a higher quality of the generated reviews. We adopt this work as our baseline and aim to develop a more computationally efficient approach than direct LLM fine-tuning while achieving higher-quality reviews. The dataset publicised by the authors of this study is being directly used by us in our work.
Ye et al. (2024) highlight key issues in LLM-generated reviews, such as giving overly high ratings in some cases and being easily influenced by authors to generate only minor limitations. Considering the contributions of this work, we reinforce the need for our approach, not as a tool for final review decisions but as a means to provide constructive feedback to researchers and reviewers.
3 Dataset
The dataset used in our study is borrowed from Yu et al. (2024), comprising papers submitted to various conferences along with multiple peer reviews per paper. Among these, only the datasets for ICLR 2024 and NeurIPS 2023 are publicly available. However, the number of GPT-4-consolidated reviews for NeurIPS 2023 is limited to just 235 samples. Hence, our analysis focuses only on the ICLR 2024 papers. As reported in Yu et al. (2024), the dataset statistics for the ICLR 2024 papers are summarized in Table 1.
Attribute Value
Total Papers 5,653
Average Tokens per Paper 9,815
Total Reviews 21,839
Acceptance Rate 37%
Domain Machine Learning
Table 1: Dataset statistics for the ICLR 2024 dataset, as reported in Yu et al. (2024).
Each paper has multiple peer reviews including sections such as Summary, Strengths, Weaknesses, Questions, Soundness, Presentation, Contribution, Confidence, and Rating to assess the quality and clarity of the paper.
From the dataset, we utilize only the Summary, Strengths, Weaknesses, and Questions sections from each review, along with the full content of the corresponding paper. Other sections, such as Ratings are excluded, as our framework is designed to provide constructive feedback rather than produce a final decision. The selected sections offer qualitative assessments of the paper’s content and provide valuable insights to guide both authors and reviewers. Additionally, we use the GPT-4 OpenAI et al. (2024) consolidated reviews dataset, sourced from Yu et al. (2024), for finetuning LLMs to generate final structured reviews. This dataset covers 20% of randomly selected ICLR-24 papers (2036 papers), where the multiple peer reviews are consolidated using GPT-4 to generate integrated reviews in a unified format.
4 Methodology
In this section, we describe the implementation of our proposed framework. Section 4.1 describes the initial passage extraction using Dense Passage Retrieval (DPR) Karpukhin et al. (2020). These extracted passages are used as training data in Section 4.3, to train a Graph Neural Network (GNN) that learns to extract relevant passages from a research paper. This trained GNN enables our model to perform competitively on the test set without the need for grounding on human-generated reviews. Post that, we describe our process of fine-tuning LLMs to generate the final reviews, in Section 4.5. An overview of our methodology can be found in Figure 2
Figure 2: The system begins with a parsed document from which DPR extracts passages for GNN training. The GNN learns to identify key passages relevant to review generation. The trained GNN selects important passages without requiring human-generated reviews. These passages, combined with a structured prompt, are input to fine-tune an LLM. The fine-tuned LLM generates the final review efficiently, without external grounding.
4.1 DPR-based Passage Extraction
Dense Passage Retrieval (DPR) is a neural retrieval model, primarily trained on question-answering datasets, that encodes both queries and context into dense vector representations. Given an input query, DPR retrieves the most relevant passages based on their semantic similarity in the dense embedding space. We use a pre-trained DPR model, to identify the most relevant passages corresponding to the different sections of the reviews that are Summary, Strengths, Weaknesses, and Questions, for each paper in the training set. Specifically, we use the context encoder [^1] https://huggingface.co/facebook/dpr-ctx_encoder-multiset-base and the query encoder [^2] https://huggingface.co/facebook/dpr-question_encoder-multiset-base from the multiset-base DPR variant.
[…]
This final set of passages represents the most relevant excerpts from the full paper that are semantically linked to the review content. These extracted passages serve as the gold-standard outputs for training the GNN, as described in Section 4.3
4.2 Graph Construction for Representing a Paper
The idea of graph construction employed in our work is inspired by Zhao et al. (2024). We borrow their idea of representing a document as a hierarchical graph but extend it beyond the sentence level. We propose to build a deeper graph, which contributes towards learning more intricate inter-dependencies of passages and sentences in a paper.
Specifically, we formalize the document structure as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents nodes and $\mathcal{E}$ denotes edges encoding hierarchical and sequential relationships.
Node Definitions: Our graph construction is designed to be generalizable across diverse academic paper structures, accommodating arbitrary levels of nesting and heterogeneous section organization.
  • Paper Node ($v_{\text{paper}}$): A single root node representing the entire document.
  • Heading Nodes $\mathcal{H}={h_{1},h_{2},\dots,h_{m}}$: Nodes corresponding to top-level section headings.
  • Sub-heading Nodes $\mathcal{S}_{i}={s_{i1},s_{i2},\dots,s_{ik}}$: Nodes representing sub-headings under each heading $h_{i}$. This structure naturally extends to accommodate multiple levels of nesting, making it suitable for documents with deeper hierarchical organization.
[…]
  • Sentence Nodes $\mathcal{T}_{jk}={t_{jk1},t_{jk2},\dots,t_{jkp}}$: Nodes for individual sentences within each passage $o_{jk}$, tokenized using NLTK Bird and Loper (2004).
Edge Definitions: Edges capture both hierarchical relationships and sequential flow within the document.
  • Root connections:
    $(v_{\text{paper}},h_{i})\in\mathcal{E}_{\text{hier}},\ \forall h_{i}\in\mathcal{H}$
  • Heading-subheading:
    $(h_{i},s{ij})\in\mathcal{E}_{\text{hier}},\ \forall s{ij}\in\mathcal{S}i$
  • Content links:
    $(s_{ij},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ if }\mathcal{S}_{i}\neq\emptyset;\\ (h_{i},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ otherwise}$
  • Passage-sentence:
    $(o_{jk},t_{jkq})\in\mathcal{E}_{\text{hier}},\ \forall t_{jkq}\in\mathcal{T}_{jk}$
  • Sequential edges (horizontal flow):
    • Passages: $(o_{jk},o_{j(k+1)})\in\mathcal{E}_{\text{seq}},\ \forall\ 1\leq k<n$
    • Sentences: $(t_{jkq},t_{jk(q+1)})\in\mathcal{E}_{\text{seq}},\ \forall\ 1\leq q<p$
The final edge set is the union of hierarchical and sequential edges:
$\mathcal{E}=\mathcal{E}_{\text{hier}}\cup\mathcal{E}_{\text{seq}}$ (1)
where $\mathcal{E}_{\text{hier}}$ captures the vertical (structural) relationships and $\mathcal{E}_{\text{seq}}$ captures the horizontal (textual progression) flow. This graph construction is agnostic to section depth and flexibly adapts to any nested structure within academic documents. A toy example of such a graph, built from a paper with 2 Headings, 2 Sub-Headings, 4 Passages, and 6 Sentences, is shown in Figure 1.
4.3 GNN Training for Passage Retrieval
To train a GNN on the graph constructed in Section 4.2, we take the data generated in Section 4.1, which maps each paper $p_{i}$ to a set of important passages, $\mathcal{O}_{i}=\{o_{1},o_{2},\dots,o_{n}\}$. Each passage is important for generating a review for the paper it is associated with. This mapping is used as a training set for training the GNN.
The node embeddings for sentence nodes are initialized with embeddings from a Sentence-Transformer Reimers and Gurevych (2019) model, namely all-mpnet-base-v2. The node embeddings for all other nodes are initialized with zero embeddings of dimension $768$, matching that of the Sentence-Transformer model.
Now, a Graph Attention Network Veličković et al. (2018) is used to train the graph to predict the important passages. After passing the initial representation of the graph nodes through three layers of GATConv [^3] https://pytorch-geometric.readthedocs.io/en/latest/generated/torch_geometric.nn.conv.GATConv.html (PyTorch Geometric’s implementation of Graph Attention Networks), we apply a non-linear classifier to the final passage node representations. The classification is defined as follows:
  • A label of 0 indicates that the passage is not important for review generation.
  • A label of 1 indicates that the passage is important for review generation.
The GNN is trained using various configurations based on the values of $k$ and $m$, as described in the DPR process in Section 4.1. These trained GNN models are denoted as GNN $(k,m)$.
We hypothesize that our choice of graph construction allows our model to capture the intricate inter-dependencies of sentences and passages across different sections within an academic paper, and effectively make a choice about which passage(s) is/are important. The dependencies, according to our hypothesis, will be learned from the sequential and hierarchical connections that we have introduced in our graph construction method. The passing of information from one node to another will be enhanced by the GATConv (Equations 2, 3, and 4) layers, which involves the concept of attention to selectively propagate messages in the Message Passing layer of a GNN.
$\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$ (2)
$e_{ij}=\text{LeakyReLU}\Bigl{(}\mathbf{a}^{T}\bigl{[}\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}\bigr{]}\Bigr{)}$ (3)
$\alpha_{ij}=\frac{\exp\bigl{(}e_{ij}\bigr{)}}{\sum_{k\in\mathcal{N}(i)}\exp\bigl{(}e_{ik}\bigr{)}}$ (4)
Additional implementation details related to this section are available in Appendix B.
Figure 3: Excerpt of a test sample where passages highlighted in cyan colour refer to relevant passages identified by GNN only and yellow colour refer to relevant passages identified by both GNN and DPR
As illustrated in an example from our test set (Figure 3), the DPR-based approach fails to retrieve a key passage outlining the main contributions in the paper, whereas GNN successfully identifies it. Thus, leveraging sequentially and hierarchically connected graph nodes and GAT in GNN proves to be more effective in identifying important passages that help to enhance the review generation process. The GNN extracted passages, along with a structured prompt, serve as the input to LLM fine-tuning, as outlined in Section 4.5.
4.4 Grounding Requirements and Inference Efficiency
The advantage of training a GNN for this important classification task is twofold. First, unlike DPR-based approaches which require grounding at inference time via access to human-generated reviews or synthesized queries, our GNN-based method operates in a fully autonomous manner. Once trained, the GNN identifies salient passages directly from the graph-structured representation of the paper, eliminating the need for any external review or human guidance during inference. This distinction is crucial for real-world deployment, where human reviews may not be available or feasible to use.
Second, by selecting only the passages deemed most relevant, we significantly reduce the input length and associated computational costs for downstream LLM-based review generation. This not only leads to higher performance of LLMs, as outlined by Hsieh et al. (2024) but also leads to higher scores of the generated reviews (compared to our baselines), as reflected in our performance metrics (see Section 5). The statistics related to the decrease in number of passages and token length post-extraction by the GNN Model are reported in Table 2.
Input to LLM Avg Token Len Avg Passages
Full Paper 9445.92 47.07
GNN $(5,5)$ 4147.15 29.49
GNN $(5,3)$ 2899.59 21.00
Table 2: Average token length (Avg. Token Len) and Average passage count (Avg. Passages) for a full paper, our best-configuration: GNN $(5,5)$, and our fastest-configuration: GNN $(5,3)$.
4.5 Fine-Tuning Large Language Models for Review Generation
General-purpose LLMs are not trained specifically to generate peer reviews for academic research papers. Our motivation for fine-tuning is to ensure that the models learn to generate qualitatively sound, factual and structured peer reviews.
We select Mistral-Instruct-v0.2 (Mistral) Jiang et al. (2023), inspired by our baseline SEA-EA, which also uses Mistral-Instruct-v0.2 as its underlying backbone model. We also use Llama 3.1 8B (Llama) Dubey et al. (2024), as it is an open-weight state-of-the-art (SOTA) model with strong text generation capabilities for a variety of tasks. We employ LoRA (Low-Rank Adaptation) Hu et al. (2022), for fine-tuning due to its efficiency in adapting large language models to particular tasks while maintaining computational feasibility. This allows us to leverage the pre-trained capabilities of Llama and Mistral while efficiently adapting them for review generation.
We use the trained GNN, as described in Section 4.3, to retrieve relevant passages for each paper that serve as the LLM’s input during fine-tuning. The GPT-4 consolidated reviews for the ICLR 2024 papers, as released by Yu et al. (2024), are regarded as gold standard references for training the LLM to generate final reviews. The dataset split is summarized in Table 3.
Dataset Split Number of Samples
Training Set 1,765
Test Set 271
Table 3: Number of samples used for training and testing sourced from Yu et al. (2024)
Both Mistral and Llama are fine-tuned on the passages retrieved, using various configurations of the trained GNN $(k,m)$, as described in Section 4.3. These fine-tuned models are denoted Mistral $(k,m)$ and Llama $(k,m)$, respectively. The prompt and hyperparameter settings used during fine-tuning are provided in Appendix C.
5 Results and Analysis
Model ROUGE F1 ROUGE Recall BERTScore-F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
SEA-E 34.29 11.71 14.98 23.28 7.90 10.13 84.04
SEA-EA 36.97 13.02 15.88 25.62 8.97 10.97 84.15
Mistral $(5,3)$ 53.44 15.47 22.39 53.18 15.42 22.31 86.44 4.34
Llama $(5,3)$ 54.24 15.91 22.59 53.75 15.77 22.41 86.59 3.60
Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
Table 4: Overall performance on the ICLR-24 reviews test dataset, comparing baselines from Yu et al. (2024) with our fine-tuned models. For both Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral), the configuration with $(5,5)$ achieves the best performance, while $(5,3)$ has the shortest training time. Bold highlights the best scores, and underlined values indicate the second-best. Please note that the GPU configurations for finetuning Llama and Mistral were different. For details, refer to Appendix C.
We evaluate the quality of the generated reviews using ROUGE-1, ROUGE-2, and ROUGE-L (both Recall and F1) Lin (2004), as well as BERTScore-F1 Zhang* et al. (2020), on the test set of GPT-4-consolidated reviews for ICLR 2024 papers. ROUGE measures similarity by analyzing n-grams, with recall indicating how well the generated reviews capture key information from the original papers and the F1 score balancing precision and recall. In contrast, BERTScore assesses semantic similarity in the embedding space, providing a deeper understanding of meaning beyond surface-level text matching.
5.1 Performance Comparison with Baselines
We compare the baselines SEA-E and SEA-EA, two automated review generation approaches, as introduced by Yu et al. (2024). As mentioned in Table 4, our results show that we consistently outperform the baselines in all metrics, with particularly significant improvements in the ROUGE metrics (Recall and F1). Our best-performing configuration, Llama $(5,5)$ (selected based on the highest BERTScore), achieves an improvement of 58.72% over the state-of-the-art (SOTA) baselines across all evaluation metrics. To ensure that these gains are not solely due to using a more capable LLM, we also evaluated our method using Mistral. The results demonstrate that even with Mistral, integrating our GNN-based passage selection leads to substantial improvements, confirming that the performance boost is primarily due to our graph-based retrieval strategy rather than the choice of LLM alone.
5.2 Performance of Our Model
Model Review Section ROUGE-F1 ROUGE-Recall BERTScore-F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
Mistral $(5,3)$ Questions 39.24 9.33 20.00 40.90 9.74 20.91 86.15 4.34
Weaknesses 40.12 9.28 19.56 38.83 8.98 18.98 86.75
Strengths 43.15 11.17 22.07 43.63 11.25 22.47 87.83
Summary 45.18 13.08 25.28 45.35 13.14 25.38 88.09
Llama $(5,3)$ Questions 40.15 9.66 20.02 39.53 9.47 19.85 86.26 3.60
Weaknesses 40.74 9.56 19.63 40.59 9.50 19.60 86.79
Strengths 42.93 11.13 22.09 43.36 11.20 22.44 87.84
Summary 45.65 13.60 25.85 46.67 13.92 26.45 88.23
Mistral $(5,5)$ Questions 39.20 9.00 19.91 40.67 9.32 20.83 86.09 6.44
Weaknesses 40.27 9.21 19.46 39.49 9.01 19.13 86.71
Strengths 43.45 11.49 22.34 44.11 11.61 22.82 87.84
Summary 46.24 14.44 26.32 46.56 14.57 26.48 88.33
Llama $(5,5)$ Questions 39.54 9.31 20.01 38.43 9.02 19.62 86.25 5.30
Weaknesses 40.72 9.66 19.79 40.72 9.59 19.93 86.77
Strengths 43.43 11.76 22.24 43.36 11.69 22.34 87.91
Summary 46.92 14.90 26.74 47.30 14.99 26.92 88.36
Table 5: The sectionwise performance across the ICLR-24 reviews test dataset after finetuning Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral). Per section of the review, the best results have been made bold, while the second-best results have been underlined. Please note that the GPU configurations for finetuning Llama and Mistral were different. For details, refer to Appendix C.
As shown in Table 5, our best-performing configuration, Llama $(5,5)$ consistently outperforms the baselines across ROUGE-F1, ROUGE-Recall, and BERTScore metrics. Llama $(5,5)$ configuration is finetuned with input passages inferred from the GNN $(5,5)$. This highlights its ability to generate reviews that closely align with gold-standard references. The integration of a Graph Attention Network (GAT) enables the model to effectively capture complex dependencies between relevant passages and sentences within the paper, offering the essential context required to generate comprehensive reviews.
Notably, even the fastest variant of our fine-tuned models, Llama $(5,3)$, which requires the shortest training time, performs competitively. Despite using a smaller number of retrieved passages, this variant shows only an average marginal decrease of 0.71% in performance metrics compared to the best-performing configuration, Llama $(5,5)$. Further, it consistently outperforms the state-of-the-art (SOTA) baselines, while reducing training time by 32.07%. This trend is consistently observed across all metrics for Mistral too. It highlights that our GNN-based approach can achieve near-optimal performance with significantly less computational overhead.
As shown in Table 5, a detailed analysis of our model’s performance across different review sections reveals that the Summary and Strengths sections achieve higher scores across ROUGE and BERTScore in our best-performing model, Llama $(5,5)$. This indicates that the model is effective at capturing and utilizing essential context from the paper to generate these key components of the review. Conversely, the Weaknesses and Questions sections receive comparatively lower scores. This indicates that identifying limitations and raising questions require a deeper semantic understanding of the paper as weaknesses may not always be explicitly stated in the passages of the paper. Overall, GNN helps to reflect on the content of the paper comprehensively to generate suitable reviews.
5.3 LLM-as-a-judge Evaluation
Review Review Quality Assessment Criteria
Reviewer Confidence Review Thoroughness Review Constructiveness Review Helpfulness
Gold 3.98 3.87 3.79 3.79
Mistral $(5,3)$ 3.50 3.35 3.44 3.32
Llama $(5,3)$ 3.61 3.45 3.47 3.39
Mistral $(5,5)$ 3.48 3.36 3.43 3.34
Llama $(5,5)$ 3.63 3.44 3.47 3.41
Table 6: The table reports the qualitative evaluation of model-generated reviews using Gemini-2.0-Flash as an LLM judge. It presents the average scores (out of 5) across four review quality criteria: Reviewer Confidence, Review Thoroughness, Review Constructiveness, and Review Helpfulness, in comparison to gold reviews on the test set. Bold highlights the model best scores, and underlined values indicate the second-best.
As mentioned by Zheng et al. (2023), the LLM-as-a-judge method offers an explainable and affordable way to approximate human preferences, which are otherwise costly to gather. We employ a large language model, Gemini-2.0-Flash [^4] https://deepmind.google/technologies/gemini/ as an automated judge to evaluate the quality of reviews generated by our proposed model as compared to the gold-standard reviews. The evaluation criteria are designed to capture key aspects of review quality and include:
  1. 1.
    Reviewer Confidence: reflects the reviewer’s depth of technical understanding and engagement with the content.
  2. 2.
    Review Thoroughness: measures how well the review covers the methodology, results, and reproducibility.
  3. 3.
    Review Constructiveness: assesses the clarity and usefulness of suggestions for improvement.
  4. 4.
    Review Helpfulness: indicates how well the review identifies important issues and explains them clearly for the authors.
The model is prompted to assign a score from 1 to 5 based on the above criteria. The evaluation prompt is detailed in Appendix D. As shown in Table 6, our best-performing model, Llama $(5,5)$, scores only 9.59% below gold reviews across all criteria. The fastest variant, Llama $(5,3)$, achieves comparable quality with just a 9.78% drop relative to gold reviews while reducing training time by 32.07%. All Llama and Mistral variants consistently score above 3.0, indicating that the reviews generated grasp key contributions and provide relevant feedback with actionable suggestions.
5.4 Ablation Studies
We perform ablation studies by systematically varying the values of $k\in\{1,3,5\}$ and $m\in\{3,5\}$ during LLM fine-tuning. The detailed results are reported in Appendix A. The results indicate that increasing $k$ and $m$ provides richer context to both Mistral and Llama to generate reviews, leading to consistent improvements across all evaluation metrics. Additionally, we observe that higher values of $m$ result in longer training times, as the GNN retrieves and passes a larger set of passages as input for finetuning the LLMs.
6 Conclusion
In this work, we introduce AutoRev, an automatic peer review generation system designed specifically for academic research papers. Our approach leverages a novel graph-based framework to represent the structure of a paper which enables the effective extraction of important passages using a Graph Neural Network. This strategy of selecting passages and reducing the input length to LLMs addresses the limitations of traditional methods, like finetuning using the entire paper as input Yu et al. (2024), which leads to sub-optimal performance of LLMs due to large input token length Hsieh et al. (2024).
Unlike DPR-based methods that require initial grounding on human written references to retrieve relevant passages, AutoRev employs a GNN that learns to identify important content directly from the paper’s graphical representation, without the need for grounding on existing reviews. This not only streamlines the process but also reduces computational overhead while enhancing review quality.
Our experiments on the ICLR 2024 dataset demonstrate that AutoRev consistently outperforms existing baselines, by 58.72% on popular evaluation metrics like ROUGE, and BERTScore.
Overall, AutoRev provides a promising solution for pre-review feedback, aiding both authors and reviewers in the academic publication process. Future work will explore extending this framework to a broader range of academic domains and investigating further improvements in graph construction and neural retrieval techniques to continue advancing the state-of-the-art in automated review generation.
Limitations
While our framework provides a robust and high-performing solution for review generation, we acknowledge that no automatic system can fully replicate human judgment in complex tasks such as peer review, as noted by Ye et al. (2024). Our goal is to assist authors in refining their papers and we aim for our system to support human reviewers. Currently, our system is primarily focused on ICLR papers due to the limited availability of public datasets. However, we believe in the generalization capabilities of our model, given the proven effectiveness of graph-based approaches in various downstream NLP tasks, as discussed in Section 1.
Ethical Considerations
The deployment of AutoRev raises key ethical concerns related to bias, misuse, and academic integrity. Since the system is trained on existing peer reviews, it may inherit and amplify biases present in those texts, such as leniency toward well-known authors or under-representation of certain research areas, potentially skewing evaluations. Moreover, there is a risk that users may over-rely on AutoRev for making or substituting final review decisions, which could undermine the depth and critical thinking central to human peer review. Additionally, the realistic quality of AI-generated reviews opens the possibility of unethical use, such as submitting AI-authored critiques as original human work. To address these issues, we advocate for responsible usage of AutoRev as a supplementary tool, accompanied by transparency in AI involvement and safeguards to preserve academic integrity.
Appendix A Ablation studies
The quantitative performance of different combinations of $k$ (number of review sentences concatenated; $k=1,3,5$) and $m$ (number of top DPR-retrieved passages used to train the GNN; $m=3,5$) for review generation is reported in Table 7 and Table 8, evaluated against the gold-standard reviews. While calculating the ROUGE metrics, we enable the use_stemmer option to ensure consistent matching of words with different morphological forms, to improve the reliability of the evaluation.
Model ROUGE-F1 ROUGE-Recall BERTScore F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
Mistral $(1,3)$ 53.54 15.81 22.68 53.26 15.74 22.59 86.51 7.60
Mistral $(1,5)$ 53.54 15.76 22.64 52.81 15.56 22.36 86.51 10.22
Mistral $(3,3)$ 53.86 15.80 22.62 53.21 15.63 22.36 86.54 6.33
Mistral $(3,5)$ 53.87 15.91 22.74 53.43 15.79 22.57 86.54 7.64
Mistral $(5,3)$ 53.44 15.47 22.39 53.18 15.42 22.31 86.44 4.34
Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
Table 7: Overall performance of Mistral across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
Model ROUGE-F1 ROUGE-Recall BERTScore F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
Llama $(1,3)$ 54.42 16.19 22.76 53.71 15.98 22.48 86.61 6.65
Llama $(1,5)$ 54.40 16.09 22.82 53.58 15.86 22.50 86.61 8.23
Llama $(3,3)$ 54.52 16.03 22.67 53.79 15.83 22.39 86.61 5.11
Llama $(3,5)$ 54.61 16.30 22.88 53.78 16.06 22.56 86.66 6.28
Llama $(5,3)$ 54.24 15.91 22.59 53.75 15.77 22.41 86.59 3.60
Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
Table 8: Overall performance of Llama across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
Appendix B GNN Implementation Details
We trained and tested our model on a single NVIDIA RTX6000 GPU with 48GB of VRAM.
B.1 Model Architecture
The model contains a Graph Neural Network (GNN) module and a passage classifier.
GNN Module
The GNN uses three graph attention (GATConv) layers. The first GATConv layer processes input features (of size 768 (embedding dimension of Sentence Encoder)) into 128-dimensional features with 4 attention heads, producing a concatenated output of $128\times 4$ dimensions. A residual connection is incorporated by applying a linear layer to project the input to $128\times 4$, which is added to the GATConv output. This is followed by an ELU activation and dropout ($p=0.2$).
The second GATConv layer maps the $128\times 4$ input to 256-dimensional features using 2 attention heads (output dimension $256\times 2$). A residual connection similarly projects the prior layer’s output to $256\times 2$, followed by ELU and dropout ($p=0.2$). The final GATConv layer upscales the $256\times 2$ features to the target output dimension (1024 in our case) without residual connections and dropout layers.
Passage Classifier
For passage nodes, we employ a feedforward classifier to map the output embeddings from the GNN to the final label space. The classifier is implemented as a multi-layer perceptron (MLP) comprising the following sequence of operations:
  • A linear transformation projecting the input to a 512-dimensional space.
  • A SiLU activation followed by dropout with a probability of $p=0.3$.
  • A second linear layer projecting to 256 dimensions.
  • Layer normalization applied to the 256-dimensional features.
  • Another SiLU activation.
  • A final linear layer projecting to the number of target classes (typically 2 for binary classification).
This architecture enables the model to learn nonlinear transformations of the graph-encoded passage representations while incorporating regularization and normalization for stable training.
B.2 Training
The model is trained for 30 epochs, with the best validation performance checkpoint retained for inference. The loss function used for training is the Cross Entropy Loss.
Appendix C Finetuning Implementation Details
We fine-tune Llama-3.1-8B using two NVIDIA L40S GPUs and Mistral using a single NVIDIA RTX 6000 Ada GPU. Both GPUs are equipped with 48GB of GDDR6 VRAM.
We apply Low-Rank Adaptation (LoRA) with a rank of 8 and a scaling factor ($\alpha=16$). LoRA is integrated into the following transformer layers: q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj. Models are fine-tuned for 3 epochs. The total training time for different configurations is reported in Table 7 and Table 8. The full set of hyperparameters used during training is listed in Table 9.
Hyperparameter Value
Batch Size 1
Gradient Accumulation Steps 1
16-bit Floating Point Precision True
Optimizer adamw_torch
Learning Rate $1\times 10^{-4}$
Max Gradient Norm 0.3
Warmup Ratio 0.03
LoRA Rank 8
LoRA scaling factor 16
LoRA Dropout 0.05
Weight Decay 0.01
LR Scheduler Type cosine
Table 9: Hyperparameters Used During Finetuning
Prompt for Finetuning
The prompt used for finetuning both Llama 3.1 8B and Mistral-Instruct-v0.2 can be found below:
Prompt Used During Finetuning Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. ### Instruction:
Generate a structured review for the research
paper passages provided below. The review should include a summary of the paper, its strengths, weaknesses, and questions for the authors. Consider that the paper is being reviewed for the ICLR conference. ### Research Paper Passages:
### Review for the paper:
[…]
The extracted relevant passages are inserted ahead of the "Research Paper Passages" heading. Similarly, the corresponding content for each review section (Summary, Strengths, Weaknesses, and Questions) is added after their respective headings in the generated review.
Appendix D LLM-as-a-judge evaluation
** Instruction ** : You are an expert meta-reviewer. Evaluate the following peer review text according to the scoring
[…]
Give rating for the categories described below on the full scale from 1(lowest rating) to 5(highest rating).
Assign scores independently per category (they may differ significantly).
**Category 1: Reviewer Confidence**
1 – Unqualified: Admits lack of expertise | Fundamental misunderstandings | Fails to engage with technical content
2 – Limited: Surface-level comprehension | Misses >2 key concepts | Confuses established methods
3 – Functional: Grasps main contributions | Identifies 1-2 technical flaws | Limited literature context
4 – Qualified: Precise technical critique | Contextualizes in recent work (>3 citations) |
Identifies methodological tradeoffs
5 – Authoritative: Demonstrates field leadership | Predicts follow-up research directions |
Reveals fundamental limitations in approach
**Category 2: Review Thoroughness**
1 – Negligent: <3 substantive comments | Skips methodology/results | Fails to address claims
2 – Patchy: Only addresses obvious aspects | 0 analysis of experiments | No ethical consideration
3 – Standard: Checks main claims | Basic methodology review | 1 reproducibility concern noted
4 – Rigorous: Line-by-line methodology check | Statistical validity analysis | 2+ improvement suggestions
5 – Forensic: Alternative approach proposals | Re-analysis of key results |
Compliance checklist (ethics, reproducibility, etc.)
**Category 3: Constructiveness**
1 – Destructive: Personal attacks | Unsubstantiated dismissal | 0 actionable feedback
2 – Demanding: Unrealistic requests ("redo all experiments") | Vague quality complaints
3 – Directive: General improvement areas | 1-2 specific examples | Mixed positive/negative tone
4 – Strategic: Priority-ranked suggestions | Alternative methodology paths | Conference-specific improvements
5 – Transformative: Step-by-step revision roadmap | Code/experiment snippets provided | Before-after examples
**Category 4: Helpfulness**
1 – Misleading: Factually incorrect suggestions | Contradicts paper’s goals | Harmful recommendations
2 – Abstract: "Improve writing" without examples | "Add experiments" without design
3 – Guided: Section-specific edits | 2-3 implementable suggestions | Partial error diagnosis
4 – Impactful: Directly address weaknesses | Templates/examples provided | Clear priority levels
5 – Revelatory: Exposes fundamental flaws | Provides benchmark comparisons | Supplies missing literature
**Scoring Methodology**
For EACH category:
1. Find the HIGHEST level (5-1) where ALL criteria are FULLY satisfied (Base Level)
2. Check criteria in the NEXT HIGHER LEVEL (Base Level + 1):
   - Calculate partial credit: (Satisfied Criteria in Higher Level / Total Criteria in Higher Level)
3. Final Score = Base Level + Partial Credit
4. Round to 2 decimal places (0.33 increments for 3-criteria levels, 0.5 for 2-criteria levels)
[…]
Respond ONLY with this JSON format:
  "review_scores": {{
    "Category 1": "",
    "Category 2": "",
    "Category 3": "",
    "Category 4": ""
Figure 4: Prompt used during evaluating the quality of generated reviews with LLM as a Judge
The prompt used for LLM-as-a-judge with Gemini-2.0-Flash is shown in Figure 4. The corresponding model-generated review is inserted under the appropriate subheadings within the prompt.
To demonstrate the effectiveness and sensitivity of our evaluation prompt, we include two manually crafted sample reviews: one representing a poorly written, low-quality review, and another simulating an average review. These examples are not generated by our system but are intentionally designed to test whether the LLM-as-a-judge can distinguish between different qualities of feedback. This validation step supports the reliability of the evaluation framework in scoring model outputs.
Example 1: Structured Critical Review (Moderate Quality)
Summary: The paper tackles a topic that lacks clear novelty, presents a poorly motivated approach, and delivers incomplete experiments with unconvincing results. The writing is disorganized, and key methodological and comparative details are missing, making it difficult to assess the contribution.
Strengths:
  • Addresses a relevant area of graph-based modeling (though similar to existing work).
  • Attempts to combine embeddings and graph structure in a single framework.
Weaknesses:
  • Motivation is vague and fails to establish why the problem matters.
  • No thorough comparison to closely related methods; literature review is superficial.
  • Methodology section is under-specified, lacking algorithmic details and hyperparameter settings.
  • Experimental results are incomplete: missing error bars, statistical analysis, and ablation studies.
  • Figures and tables are low resolution or poorly labeled, hindering interpretability.
Questions:
  • What is the core novelty compared to the most recent graph-embedding methods?
  • How were key hyperparameters chosen, and how sensitive are the results to those choices?
  • Can you provide ablation studies showing the contribution of each component in the model?
  • Why are error bars and statistical significance tests omitted from the reported results?
  • How does your method scale to larger datasets, and what is its computational complexity?
The evaluation scores obtained for Example-1 are as follows,
  1. 1.
    Reviewer Confidence : 3.0
  2. 2.
    Review Thoroughness : 3.0
  3. 3.
    Review Constructiveness : 3.33
  4. 4.
    Review Helpfulness : 3.0
Example 2: Informal and Unstructured Review (Low Quality)
This paper is pretty weak. The writing is sloppy and hard to follow. I’m not convinced the idea is novel. There are lots of papers out there doing similar stuff. The experiments seem incomplete, and the results aren’t very impressive. I don’t think this will be of interest to ICLR readers. You should do more experiments or maybe scrap the whole thing and start over.
Comments:
  • The motivation is unclear.
  • There’s no thorough comparison to related work.
  • The methodology description is too brief.
  • Results tables have no error bars.
  • Figures are low resolution and hard to read.
Recommendation: Reject.
The evaluation scores obtained for Example-2 are as follows,
  1. 1.
    Reviewer Confidence : 2.0
  2. 2.
    Review Thoroughness: 1.0
  3. 3.
    Review Constructiveness : 2.0
  4. 4.
    Review Helpfulness : 2.0

### v2

@@
-AutoRev: Automatic Peer Review System for Academic Research Papers
-Abstract
-Generating a review for an academic research paper is a complex task that requires a deep understanding of the document’s content and the interdependencies between its sections.
-It demands not only insight into technical details but also an appreciation of the paper’s overall coherence and structure.
-Recent methods have predominantly focused on fine-tuning large language models (LLMs) to address this challenge.
-However, they often overlook the computational and performance limitations imposed by long input token lengths.
-To address this, we introduce AutoRev, an Automatic Peer Review System for Academic Research Papers.
-Our novel framework represents an academic document as a graph, enabling the extraction of the most critical passages that contribute significantly to the review.
-This graph-based approach demonstrates effectiveness for review generation and is potentially adaptable to various downstream tasks, such as question answering, summarization, and document representation.
-When applied to review generation, our method outperforms SOTA baselines by an average of 58.72% across all evaluation metrics.
-We hope that our work will stimulate further research in applying graph-based extraction techniques to other downstream tasks in NLP.
-We plan to make our code public upon acceptance.
-AutoRev: Automatic Peer Review System for Academic Research Papers
-Maitreya Prafulla Chitale1thanks: Equal contribution.
-Ketaki Mangesh Shetye [^1] footnotemark: 1 Harshit Gupta1Manav Chaudhary1Vasudeva Varma1 1IIIT Hyderabad {maitreya.chitale, ketaki.shetye, harshit.g, manav.chaudhary}@research.iiit.ac.in vv@iiit.ac.in
-1 Introduction
-Figure 1: Example graph constructed out of a paper consisting of 2 Headings, 2 Sub-Headings, 4 Passages, and 6 Sentences.
-Recent research has increasingly focused on developing automatic systems for peer review, including tools that assist reviewers, assess peer-review processes, and directly generate reviews for academic papers Yu et al.
-(2024); Aitymbetov and Zorbas (2025); Church et al.
-(2025).
-Concurrently, there has been growing interest in examining the strengths and limitations of such systems, particularly the use of large language models (LLMs) for automated review generation Ye et al.
-(2024).
+AutoRev: Multi-Modal Graph Retrieval for Automated Peer-Review Generation
+Abstract.
+Enhancing the quality and efficiency of academic publishing is critical for both authors and reviewers, as research papers are central to scholarly communication and a major source of high-quality content on the web.
+To support this goal, we propose AutoRev, an automatic peer-review system designed to provide actionable, high-quality feedback to both reviewers and authors.
+AutoRev leverages a novel Multi-Modal Retrieval-Augmented Generation (RAG) framework that combines textual and graphical representations of academic papers.
+By modelling documents as graphs, AutoRev effectively retrieves the most pertinent information, significantly reducing the input context length for LLMs and thereby enhancing their review generation capabilities.
+Experimental results show that AutoRev outperforms state-of-the-art baselines by up to 58.72% and demonstrates competitive performance in human evaluations against ground truth reviews.
+We envision AutoRev as a powerful tool to streamline the peer-review workflow, alleviating challenges and enabling scalable, high-quality scholarly publishing.
+By guiding both authors and reviewers, AutoRev has the potential to accelerate the dissemination of quality research on the web at a larger scale.
+Code will be released upon acceptance.
+Graph Neural Networks (GNNs), Retrieval Augmented Generation (RAG), Large Language Models (LLMs), Automatic Peer Reviewing
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Information extractionccs: Computing methodologies Natural language generationccs: Information systems Language modelsccs: Information systems Document structure
+1. Introduction
+Figure 1. Example document graph with nodes for headings, sub-headings, passages, and sentences. This structure allows AutoRev to capture hierarchical and sequential dependencies across sections.:
+Recent research has increasingly focused on developing automatic systems for peer review, including tools that assist reviewers, assess peer-review processes, and directly generate reviews for academic papers (Yu et al., 2024; Aitymbetov and Zorbas, 2025; D’Arcy et al., 2024; Idahl and Ahmadi, 2025).
+Concurrently, there has been growing interest in examining the strengths and limitations of such systems, particularly the use of large language models (LLMs) for automated review generation (Ye et al., 2024).
 Motivated by both of these trends, we propose AutoRev, a system designed to generate peer reviews from a guidance perspective, offering constructive feedback to support researchers and reviewers.
-With the rise of Graph Neural Networks, the use of graphs as extractors or as guides for facilitating downstream tasks has recently gained traction in the NLP community Chitale et al.
-(2025); Gao et al.
-(2025); Agarwal et al.
-(2025).
-Inspired by this, we propose to use graphs in the domain of academic paper analysis, specifically, generating reviews that closely replicate those produced by existing peer review systems in conferences.
-AutoRev can generate high-quality reviews without relying on any prior human evaluations of the paper.
+With the rise of Graph Neural Networks, the use of graphs as extractors or as guides for facilitating downstream tasks has recently gained traction in the NLP community (Chitale et al., 2025; Gao et al., 2025; Agarwal et al., 2025; Zhu et al., 2025).
+Inspired by this, we propose to use graphs in the domain of academic paper analysis, specifically, in generating reviews that closely replicate the structure and quality of those produced by existing peer review systems in conferences.
 Furthermore, by leveraging the graph’s capability as an extractor of passages, we significantly reduce the input token length to the LLM.
-Reducing the token length of the input has been shown to enhance LLM performance Hsieh et al.
-(2024).
+Reducing the token length of the input is shown to enhance LLM performance (Hsieh et al., 2024).
 To summarize our main contributions:
+  1. (1)
     We propose a system for generating concise, structured reviews (Summary, Strengths, Weaknesses, and Questions) to assist authors and reviewers.
-  • To our knowledge, this is the first approach using graphs in this domain.
-  • Our approach improves over baselines by 58.72% on average across all metrics, without relying on human-written reviews and staying efficient.
-It is important to highlight that this paper does not seek to determine the final reviews of a submission.
+To our knowledge, this is the first approach which uses graphs in this domain.
+  2. (2)
+    Our approach improves over SoTA baselines by 58.72% on average across all metrics while being computationally efficient.
+  3. (3)
+    We conduct thorough human evaluations and other automatic evaluations to demonstrate the competitive performance of our method against ground truth reviews.
+It is important to highlight that this paper does not seek to determine the final reviews of a paper.
 Rather, we present our framework as a tool to provide timely feedback to researchers and serve as a valuable guide for reviewers.
-2 Related Works
+We envision AutoRev as a step towards scalable, automated peer-review systems that can complement human judgment.
+With the continued surge in conference submissions across disciplines, the demand for such tools is increasingly pressing [^1] https://aaai.org/aaai-launches-ai-powered-peer-review-assessment-system/.
+By alleviating the bottlenecks of the traditional peer-review process, AutoRev can help accelerate the dissemination of high-quality research on the web, ensuring that credible scholarly knowledge remains accessible at scale.
+2. Related Works
 A significant amount of research has been conducted on aiding peer reviews for academic papers.
-Some studies explore the involvement of AI in the selection process Aitymbetov and Zorbas (2025), while some studies work on generating meta-reviews Kumar et al.
-(2023).
-Sufficient discussion and avenues have been explored where AI/NLP techniques can contribute towards aiding automatic peer reviews Kuznetsov et al.
-(2024).
-Though, to the best of our knowledge, significant research has not been done in the field of generating a whole review for an academic paper.
-A recent study by Faizullah et al.
-(2024) explores a DPR-based approach for generating suggestive limitations.
-It highlights the necessity of using DPR to address length constraints and extract only the most relevant passages from a paper.
-This study served as an initial motivation for incorporating DPR into our approach.
-Another notable work by Yu et al.
-(2024) introduces an automated paper reviewing framework.
-The work also includes a standardized dataset of GPT-4 OpenAI et al.
-(2024) consolidated reviews collected from various academic venues.
-To enhance review generation, the authors propose a self-correction strategy during LLM fine-tuning to ensure a higher quality of the generated reviews.
-We adopt this work as our baseline and aim to develop a more computationally efficient approach than direct LLM fine-tuning while achieving higher-quality reviews.
-The dataset publicised by the authors of this study is being directly used by us in our work.
+Some studies explore the involvement of AI in the reviewer selection process (Aitymbetov and Zorbas, 2025), while others focus on generating meta-reviews (Kumar et al., 2023).
+Prior work has also examined how AI/NLP techniques can contribute to various aspects of the peer-review pipeline (Kuznetsov et al., 2024).
+Faizullah et al.
+(2024) explore a DPR-based approach for generating suggestive limitations, highlighting the necessity of retrieval methods to overcome length constraints.
+Building on this, we use DPR outputs as the training data for our Graph Neural Network (GNN), enabling it to learn to identify salient passages for review generation.
+Yu et al.
+(2024) introduce an automated paper reviewing framework (SEA) that standardizes reviews using GPT-4222https://arxiv.org/abs/2303.08774 consolidation and proposes a self-correction strategy during fine-tuning.
 Ye et al.
-(2024) highlight key issues in LLM-generated reviews, such as giving overly high ratings in some cases and being easily influenced by authors to generate only minor limitations.
-Considering the contributions of this work, we reinforce the need for our approach, not as a tool for final review decisions but as a means to provide constructive feedback to researchers and reviewers.
-3 Dataset
+(2024) further caution that LLM-generated reviews may overrate papers or produce superficial feedback, reinforcing our design goal of providing a structured and constructive review rather than final accept/reject decisions.
+Among the most relevant recent systems, MARG (D’Arcy et al., 2024) employs a multi-agent framework, distributing different sections of a paper to specialized GPT-4 agents.
+This enables the model to handle long papers and produce more specific comments, but it comes at the cost of heavy computation and frequent inter-agent coordination errors.
+OpenReviewer (Idahl and Ahmadi, 2025) instead fully fine-tunes Llama-3.1-8B-Instruct on 79k reviews from ICLR and NeurIPS to generate structured reviews, adhering to the conference guidelines.
+This method of large-scale full fine-tuning is very costly.
+We adopt SEA (Yu et al., 2024) as our primary baseline since it is closest in scope to our work, representing a recent state-of-the-art framework for automated peer review generation on the same ICLR dataset.
+Other systems such as MARG and OpenReviewer differ substantially in design.
+MARG uses multi-agent GPT-4 orchestration, while OpenReviewer relies on large-scale fine-tuning, making direct benchmarking less meaningful.
+Thus, we benchmark against SEA for a fair and reproducible comparison.
+The integration of graph-based techniques into Retrieval Augmented Generation (RAG) (Lewis et al., 2020) is an emerging research area that aims to enhance the reasoning and factual grounding of LLMs.
+A recent comprehensive survey by Zhu et al.
+(2025) provides a foundational taxonomy for this field, categorizing systems based on their database construction, retrieval algorithms, and processing pipelines.
+Within this framework, AutoRev is a domain-specific Graph-RAG application that generates a hierarchical graph directly from a scientific paper’s text.
+It employs a learning-based Graph Attention Network (GAT) (Veličković et al., 2018) for retrieval that reduces the context length of input to LLMs, reducing the computational costs.
+3. Dataset
 The dataset used in our study is borrowed from Yu et al.
 (2024), comprising papers submitted to various conferences along with multiple peer reviews per paper.
-Among these, only the datasets for ICLR 2024 and NeurIPS 2023 are publicly available.
-However, the number of GPT-4-consolidated reviews for NeurIPS 2023 is limited to just 235 samples.
+Among these, the ICLR 2024 and NeurIPS 2023 datasets are publicly released, though the GPT-4–consolidated reviews for NeurIPS 2023 are not released.
 Hence, our analysis focuses only on the ICLR 2024 papers.
@@
 (2024), the dataset statistics for the ICLR 2024 papers are summarized in Table 1.
+Table 1. Dataset statistics for the ICLR 2024 dataset, as reported in Yu et al. (2024).
 Attribute Value
@@
 Domain Machine Learning
-Table 1: Dataset statistics for the ICLR 2024 dataset, as reported in Yu et al. (2024).
 Each paper has multiple peer reviews including sections such as Summary, Strengths, Weaknesses, Questions, Soundness, Presentation, Contribution, Confidence, and Rating to assess the quality and clarity of the paper.
@@
 The selected sections offer qualitative assessments of the paper’s content and provide valuable insights to guide both authors and reviewers.
-Additionally, we use the GPT-4 OpenAI et al.
-(2024) consolidated reviews dataset, sourced from Yu et al.
-(2024), for finetuning LLMs to generate final structured reviews.
-This dataset covers 20% of randomly selected ICLR-24 papers (2036 papers), where the multiple peer reviews are consolidated using GPT-4 to generate integrated reviews in a unified format.
-4 Methodology
-In this section, we describe the implementation of our proposed framework.
-Section 4.1 describes the initial passage extraction using Dense Passage Retrieval (DPR) Karpukhin et al.
+For our finetuning process, we utilize the GPT-4 consolidated reviews dataset introduced by Yu et al.
 (2024).
-These extracted passages are used as training data in Section 4.3, to train a Graph Neural Network (GNN) that learns to extract relevant passages from a research paper.
-This trained GNN enables our model to perform competitively on the test set without the need for grounding on human-generated reviews.
-Post that, we describe our process of fine-tuning LLMs to generate the final reviews, in Section 4.5.
-An overview of our methodology can be found in Figure 2
-Figure 2: The system begins with a parsed document from which DPR extracts passages for GNN training.
+This dataset is constructed by using GPT-4 to consolidate multiple human-written peer reviews from OpenReview [^3] https://openreview.net/ into a single, unified review for each paper.
+4. Methodology
+Figure 2.
+The system begins with a parsed document from which DPR extracts passages for GNN training.
 The GNN learns to identify key passages relevant to review generation.
@@
 These passages, combined with a structured prompt, are input to fine-tune an LLM.
-The fine-tuned LLM generates the final review efficiently, without external grounding.
-4.1 DPR-based Passage Extraction
+The fine-tuned LLM generates the final review efficiently, without external grounding.:
+In this section, we describe the implementation of our proposed framework.
+Section 4.1 describes the initial passage extraction using Dense Passage Retrieval (DPR) (Karpukhin et al., 2020).
+The extracted passages are used as training data in Section 4.3, to train a Graph Neural Network (GNN) that learns to extract relevant passages from a research paper.
+The trained GNN enables our model to perform competitively on the test set without the need for grounding on human-generated reviews, as discussed in Section 4.4.
+Post that, we describe our process of fine-tuning LLMs to generate the final reviews, in Section 4.5.
+An overview of our methodology can be found in Figure 2.
+4.1. DPR-based Passage Extraction
 Dense Passage Retrieval (DPR) is a neural retrieval model, primarily trained on question-answering datasets, that encodes both queries and context into dense vector representations.
@@
 This final set of passages represents the most relevant excerpts from the full paper that are semantically linked to the review content.
-These extracted passages serve as the gold-standard outputs for training the GNN, as described in Section 4.3
-4.2 Graph Construction for Representing a Paper
+These extracted passages serve as the ground truth labels for training the GNN, as described in Section 4.3
+4.2. Graph Construction for Representing a Paper
 The idea of graph construction employed in our work is inspired by Zhao et al.
 (2024).
-We borrow their idea of representing a document as a hierarchical graph but extend it beyond the sentence level.
+We borrow their idea of representing a document as a hierarchical graph but extend it beyond sentence level.
 We propose to build a deeper graph, which contributes towards learning more intricate inter-dependencies of passages and sentences in a paper.
-Specifically, we formalize the document structure as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents nodes and $\mathcal{E}$ denotes edges encoding hierarchical and sequential relationships.
-Node Definitions: Our graph construction is designed to be generalizable across diverse academic paper structures, accommodating arbitrary levels of nesting and heterogeneous section organization.
+Specifically, we formalize the document structure as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where nodes $\mathcal{V}$ represent text segments with semantically encoded features, and $\mathcal{E}$ denote the hierarchical and sequential relationships that form the document’s structure.
+Our graph construction is designed to be generalizable across diverse academic paper structures, accommodating arbitrary levels of nesting and heterogeneous section organization.
+4.2.1. Node Definitions
   • Paper Node ($v_{\text{paper}}$): A single root node representing the entire document.
-  • Heading Nodes $\mathcal{H}={h_{1},h_{2},\dots,h_{m}}$: Nodes corresponding to top-level section headings.
-  • Sub-heading Nodes $\mathcal{S}_{i}={s_{i1},s_{i2},\dots,s_{ik}}$: Nodes representing sub-headings under each heading $h_{i}$.
+  • Heading Nodes ($\mathcal{H}=\{h_{1},h_{2},\dots,h_{m}\}$): Nodes corresponding to top-level section headings.
+  • Sub-heading Nodes ($\mathcal{S}_{i}=\{s_{i1},s_{i2},\dots,s_{ik}\}$): Nodes representing sub-headings under each heading $h_{i}$.
 This structure naturally extends to accommodate multiple levels of nesting, making it suitable for documents with deeper hierarchical organization.
 […]
-  • Sentence Nodes $\mathcal{T}_{jk}={t_{jk1},t_{jk2},\dots,t_{jkp}}$: Nodes for individual sentences within each passage $o_{jk}$, tokenized using NLTK Bird and Loper (2004).
-Edge Definitions: Edges capture both hierarchical relationships and sequential flow within the document.
-  • Root connections:
-    $(v_{\text{paper}},h_{i})\in\mathcal{E}_{\text{hier}},\ \forall h_{i}\in\mathcal{H}$
-  • Heading-subheading:
-    $(h_{i},s{ij})\in\mathcal{E}_{\text{hier}},\ \forall s{ij}\in\mathcal{S}i$
+  • Sentence Nodes ($\mathcal{T}_{jk}=\{t_{jk1},t_{jk2},\dots,t_{jkp}\}$): Nodes for individual sentences within each passage $o_{jk}$, tokenized using NLTK (Bird and Loper, 2004).
+4.2.2. Edge Definitions
+Edges capture both hierarchical and sequential relationships within the document.
+  • Root Connections: $(v_{\text{paper}},h_{i})\in\mathcal{E}_{\text{hier}}$ for all $h_{i}\in\mathcal{H}$.
+  • Heading-SubHeading Connections:
+    $(h_{i},s_{ij})\in\mathcal{E}_{\text{hier}}$ for all $s_{ij}\in\mathcal{S}_{i}$.
   • Content Links:
-    $(s_{ij},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ if }\mathcal{S}_{i}\neq\emptyset;\\ (h_{i},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ otherwise}$
-  • Passage-sentence:
-    $(o_{jk},t_{jkq})\in\mathcal{E}_{\text{hier}},\ \forall t_{jkq}\in\mathcal{T}_{jk}$
-  • Sequential edges (horizontal flow):
-    • Passages: $(o_{jk},o_{j(k+1)})\in\mathcal{E}_{\text{seq}},\ \forall\ 1\leq k<n$
-    • Sentences: $(t_{jkq},t_{jk(q+1)})\in\mathcal{E}_{\text{seq}},\ \forall\ 1\leq q<p$
+    $(s_{ij},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ if }\mathcal{S}_{i}\neq\emptyset;(h_{i},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ otherwise.}$
+  • Passage-Sentence Connections:
+    $(o_{jk},t_{jkq})\in\mathcal{E}_{\text{hier}}$ for all $t_{jkq}\in\mathcal{T}_{jk}$.
+  • Sequential Edges:
+    For passages: $(o_{jk},o_{j(k+1)})\in\mathcal{E}_{\text{seq}}$ for $1\leq k<n$.
+    For sentences: $(t_{jkq},t_{jk(q+1)})\in\mathcal{E}_{\text{seq}}$ for $1\leq q<p$.
 The final edge set is the union of hierarchical and sequential edges:
-$\mathcal{E}=\mathcal{E}_{\text{hier}}\cup\mathcal{E}_{\text{seq}}$ (1)
+(1) $\mathcal{E}=\mathcal{E}_{\text{hier}}\cup\mathcal{E}_{\text{seq}},$
 where $\mathcal{E}_{\text{hier}}$ captures the vertical (structural) relationships and $\mathcal{E}_{\text{seq}}$ captures the horizontal (textual progression) flow.
@@
 A toy example of such a graph, built from a paper with 2 Headings, 2 Sub-Headings, 4 Passages, and 6 Sentences, is shown in Figure 1.
-4.3 GNN Training for Passage Retrieval
+4.3. GNN Training for Multi Modal Passage Retrieval
+Figure 3.
+Excerpt of a test sample where passages highlighted in cyan colour refer to relevant passages identified by GNN only and yellow colour refer to relevant passages identified by both GNN and DPR:
 To train a GNN on the graph constructed in Section 4.2, we take the data generated in Section 4.1, which maps each paper $p_{i}$ to a set of important passages, $\mathcal{O}_{i}=\{o_{1},o_{2},\dots,o_{n}\}$.
@@
 This mapping is used as a training set for training the GNN.
-The node embeddings for sentence nodes are initialized with embeddings from a Sentence-Transformer Reimers and Gurevych (2019) model, namely all-mpnet-base-v2.
+The node embeddings for sentence nodes are initialized with embeddings from a Sentence-Transformer (Reimers and Gurevych, 2019) model, namely all-mpnet-base-v2.
 The node embeddings for all other nodes are initialized with zero embeddings of dimension $768$, matching that of the Sentence-Transformer model.
-Now, a Graph Attention Network Veličković et al.
-(2018) is used to train the graph to predict the important passages.
-After passing the initial representation of the graph nodes through three layers of GATConv [^3] https://pytorch-geometric.readthedocs.io/en/latest/generated/torch_geometric.nn.conv.GATConv.html (PyTorch Geometric’s implementation of Graph Attention Networks), we apply a non-linear classifier to the final passage node representations.
+Now, a Graph Attention Network (Veličković et al., 2018) is used to train the graph to predict the important passages.
+After passing the initial representation of the graph nodes through three layers of GATConv (PyTorch Geometric’s implementation of Graph Attention Networks) [^6] https://pytorch-geometric.readthedocs.io/en/latest/generated/torch_geometric.nn.conv.GATConv.html, we apply a non-linear classifier to the final passage node representations.
 The classification is defined as follows:
+  1. (1)
     A label of 0 indicates that the passage is not important for review generation.
+  2. (2)
     A label of 1 indicates that the passage is important for review generation.
 The GNN is trained using various configurations based on the values of $k$ and $m$, as described in the DPR process in Section 4.1. These trained GNN models are denoted as GNN $(k,m)$.
-We hypothesize that our choice of graph construction allows our model to capture the intricate inter-dependencies of sentences and passages across different sections within an academic paper, and effectively make a choice about which passage(s) is/are important.
-The dependencies, according to our hypothesis, will be learned from the sequential and hierarchical connections that we have introduced in our graph construction method.
-The passing of information from one node to another will be enhanced by the GATConv (Equations 2, 3, and 4) layers, which involves the concept of attention to selectively propagate messages in the Message Passing layer of a GNN.
-$\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$ (2)
-$e_{ij}=\text{LeakyReLU}\Bigl{(}\mathbf{a}^{T}\bigl{[}\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}\bigr{]}\Bigr{)}$ (3)
-$\alpha_{ij}=\frac{\exp\bigl{(}e_{ij}\bigr{)}}{\sum_{k\in\mathcal{N}(i)}\exp\bigl{(}e_{ik}\bigr{)}}$ (4)
+We hypothesize that our choice of graph construction which models the semantic representation of nodes and flow of information through structure, allows our model to capture the intricate inter-dependencies of sentences and passages across different sections within an academic paper, and effectively make a choice about which passage(s) is/are important.
+The dependencies, according to our hypothesis, are learned from the sequential and hierarchical connections that we introduce in our graph construction method.
+The passing of information from one node to another is enhanced by the GATConv (Equations 2, 3, and 4) layers, which involves the concept of attention to selectively propagate messages in the Message Passing layer of a GNN.
+(2) $\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$
+(3) $e_{ij}=\text{LeakyReLU}\Bigl(\mathbf{a}^{T}\bigl[\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}\bigr]\Bigr)$
+(4) $\alpha_{ij}=\frac{\exp\bigl(e_{ij}\bigr)}{\sum_{k\in\mathcal{N}(i)}\exp\bigl(e_{ik}\bigr)}$
 Additional implementation details related to this section are available in Appendix B.
-Figure 3: Excerpt of a test sample where passages highlighted in cyan colour refer to relevant passages identified by GNN only and yellow colour refer to relevant passages identified by both GNN and DPR
-As illustrated in an example from our test set (Figure 3), the DPR-based approach fails to retrieve a key passage outlining the main contributions in the paper, whereas GNN successfully identifies it.
-Thus, leveraging sequentially and hierarchically connected graph nodes and GAT in GNN proves to be more effective in identifying important passages that help to enhance the review generation process.
-The GNN extracted passages, along with a structured prompt, serve as the input to LLM fine-tuning, as outlined in Section 4.5.
-4.4 Grounding Requirements and Inference Efficiency
-The advantage of training a GNN for this important classification task is twofold.
-First, unlike DPR-based approaches which require grounding at inference time via access to human-generated reviews or synthesized queries, our GNN-based method operates in a fully autonomous manner.
-Once trained, the GNN identifies salient passages directly from the graph-structured representation of the paper, eliminating the need for any external review or human guidance during inference.
-This distinction is crucial for real-world deployment, where human reviews may not be available or feasible to use.
+As illustrated in the example from our test set (Figure 3), the DPR-based approach fails to retrieve a key passage related to the main contributions of the paper, whereas our GNN successfully identifies it.
+Therefore, our multi-modal graph, which leverages both sequential and hierarchical node connections, more effectively identifies the salient passages required to enhance the review generation process.
+The GNN extracted passages, along with a structured prompt, serve as the input in the LLM fine-tuning stage, as discussed in Section 4.5.
+4.4. Grounding Requirements and Inference Efficiency
+For practical deployment, our GNN-based method offers a crucial advantage by operating autonomously at inference time.
+It identifies important passages directly from the paper’s graph structure, eliminating the need for grounding from human-generated reviews which act as queries in the DPR-based approaches, but may be unavailable in real-world scenarios.
 Second, by selecting only the passages deemed most relevant, we significantly reduce the input length and associated computational costs for downstream LLM-based review generation.
 This not only leads to higher performance of LLMs, as outlined by Hsieh et al.
-(2024) but also leads to higher scores of the generated reviews (compared to our baselines), as reflected in our performance metrics (see Section 5).
+(2024) but also leads to higher scores of the generated reviews compared to our baselines (see Section 5).
 The statistics related to the decrease in number of passages and token length post-extraction by the GNN model are reported in Table 2.
+Table 2. Average token length (Avg. Token Len) and Average passage count (Avg. Passages) for a full paper, our best-configuration: GNN $(5,5)$, and our fastest-configuration: GNN $(5,3)$.
 Input to LLM Avg Token Len Avg Passages
@@
 GNN $(5,3)$ 2899.59 21.00
-Table 2: Average token length (Avg. Token Len) and Average passage count (Avg. Passages) for a full paper, our best-configuration: GNN $(5,5)$, and our fastest-configuration: GNN $(5,3)$.
-4.5 Fine-Tuning Large Language Models for Review Generation
+4.5. Fine-Tuning Large Language Models for Review Generation
 General-purpose LLMs are not trained specifically to generate peer reviews for academic research papers.
 Our motivation for fine-tuning is to ensure that the models learn to generate qualitatively sound, factual and structured peer reviews.
-We select Mistral-Instruct-v0.2 (Mistral) Jiang et al.
-(2023), inspired by our baseline SEA-EA, which also uses Mistral-Instruct-v0.2 as its underlying backbone model.
-We also use Llama 3.1 8B (Llama) Dubey et al.
-(2024), as it is an open-weight state-of-the-art (SOTA) model with strong text generation capabilities for a variety of tasks.
-We employ LoRA (Low-Rank Adaptation) Hu et al.
-(2022), for fine-tuning due to its efficiency in adapting large language models to particular tasks while maintaining computational feasibility.
+We select Mistral-Instruct-v0.2 (Mistral) [^7] https://arxiv.org/abs/2310.06825, inspired by our baseline SEA-EA, which also uses Mistral-Instruct-v0.2 as its underlying backbone model.
+We also use Llama 3.1 8B (Llama) [^8] https://arxiv.org/abs/2407.21783, as it is an open-weight state-of-the-art (SOTA) model with strong text generation capabilities for a variety of tasks.
+We employ LoRA (Low-Rank Adaptation) (Hu et al., 2022), for fine-tuning due to its efficiency in adapting large language models to particular tasks while maintaining computational feasibility.
 This allows us to leverage the pre-trained capabilities of Llama and Mistral while efficiently adapting them for review generation.
 We use the trained GNN, as described in Section 4.3, to retrieve relevant passages for each paper that serve as the LLM’s input during fine-tuning.
-The GPT-4 consolidated reviews for the ICLR 2024 papers, as released by Yu et al.
-(2024), are regarded as gold standard references for training the LLM to generate final reviews.
+The human-written peer reviews for ICLR 2024 papers, which are consolidated with GPT-4, served as the ground truth for fine-tuning the LLM used to generate the final reviews.
+These reviews are released as a part of the dataset released by Yu et al.
+(2024).
 The dataset split is summarized in Table 3.
+Table 3. Number of samples used for training and testing sourced from Yu et al. (2024)
 Dataset Split Number of Samples
@@
 Test Set 271
-Table 3: Number of samples used for training and testing sourced from Yu et al. (2024)
 Both Mistral and Llama are fine-tuned on the passages retrieved, using various configurations of the trained GNN $(k,m)$, as described in Section 4.3.
@@
 The prompt and hyperparameter settings used during fine-tuning are provided in Appendix C.
-5 Results and Analysis
+5. Results and Analysis
+Table 4.
+Overall performance on the ICLR-24 reviews test dataset, comparing M-7B, SEA-E, SEA-EA as baselines from Yu et al.
+(2024) with our fine-tuned models.
+For both Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral), the configuration with $(5,5)$ achieves the best performance, while $(5,3)$ has the shortest training time.
+Bold highlights the best scores, and underlined values indicate the second-best.
+Please note that the GPU configurations for finetuning Llama and Mistral were different.
+For details, refer to Appendix C.
 Model ROUGE F1 ROUGE Recall BERTScore-F1 Training Time (hrs)
 R-1 R-2 R-L R-1 R-2 R-L
+M-7B 22.55 8.89 11.28 13.48 5.29 16.73 83.79
 SEA-E 34.29 11.71 14.98 23.28 7.90 10.13 84.04
@@
 Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
-Table 4: Overall performance on the ICLR-24 reviews test dataset, comparing baselines from Yu et al.
-(2024) with our fine-tuned models.
-For both Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral), the configuration with $(5,5)$ achieves the best performance, while $(5,3)$ has the shortest training time.
-Bold highlights the best scores, and underlined values indicate the second-best.
-Please note that the GPU configurations for finetuning Llama and Mistral were different.
-For details, refer to Appendix C.
-We evaluate the quality of the generated reviews using ROUGE-1, ROUGE-2, and ROUGE-L (both Recall and F1) Lin (2004), as well as BERTScore-F1 Zhang* et al.
-(2020), on the test set of GPT-4-consolidated reviews for ICLR 2024 papers.
-ROUGE measures similarity by analyzing n-grams, with recall indicating how well the generated reviews capture key information from the original papers and the F1 score balancing precision and recall.
-In contrast, BERTScore assesses semantic similarity in the embedding space, providing a deeper understanding of meaning beyond surface-level text matching.
-5.1 Performance Comparison with Baselines
-We compare the baselines SEA-E and SEA-EA, two automated review generation approaches, as introduced by Yu et al.
+This section details our system’s performance analysis.
+Section 5.1 reports the evaluation of generated review quality using ROUGE-1, ROUGE-2, and ROUGE-L (Recall and F1) (Lin, 2004), as well as BERTScore-F1 (Zhang* et al., 2020), on the GPT-4-consolidated reviews test set for ICLR 2024 papers.
+ROUGE measures n-gram overlap, with recall reflecting content coverage and F1 balancing precision and recall, while BERTScore captures semantic similarity in embedding space.
+Section 5.2 further employs the Aspect Coverage (ACov) metric (Yuan et al., 2022), designed specifically for assessing peer-review quality.
+Given the limitations of automated metrics, we also conduct human evaluations (Section 5.3) to assess review acceptability.
+To address the inherent scalability challenges of manual assessment, we employ an LLM-as-a-Judge framework (Section 5.4), which shows strong correlation with human judgments.
+5.1. Performance Comparison with Baselines
+We compare the performance of our models against baselines sourced from Yu et al.
 (2024).
-As mentioned in Table 4, our results show that we consistently outperform the baselines in all metrics, with particularly significant improvements in the ROUGE metrics (Recall and F1).
+As reported in Table 4, the M-7B baseline performs direct inference with Mistral guided by an instruction prompt to generate peer reviews, SEA-E baseline leverages standardized fine-tuning data to produce more constructive reviews, while the SEA-EA baseline further introduces a mismatch score to evaluate the consistency between paper content and generated reviews.
+Our results suggest that we consistently outperform the baselines on all metrics, with significant improvements in the ROUGE metrics (Recall and F1).
 Our best-performing configuration, Llama $(5,5)$ (selected based on the highest BERTScore), achieves an improvement of 58.72% over the state-of-the-art (SOTA) baselines across all evaluation metrics.
-To ensure that these gains are not solely due to using a more capable LLM, we also evaluated our method using Mistral.
-The results demonstrate that even with Mistral, integrating our GNN-based passage selection leads to substantial improvements, confirming that the performance boost is primarily due to our graph-based retrieval strategy rather than the choice of LLM alone.
-5.2 Performance of Our Model
-Model Review Section ROUGE-F1 ROUGE-Recall BERTScore-F1 Training Time (hrs)
-R-1 R-2 R-L R-1 R-2 R-L
-Mistral $(5,3)$ Questions 39.24 9.33 20.00 40.90 9.74 20.91 86.15 4.34
-Weaknesses 40.12 9.28 19.56 38.83 8.98 18.98 86.75
-Strengths 43.15 11.17 22.07 43.63 11.25 22.47 87.83
-Summary 45.18 13.08 25.28 45.35 13.14 25.38 88.09
-Llama $(5,3)$ Questions 40.15 9.66 20.02 39.53 9.47 19.85 86.26 3.60
-Weaknesses 40.74 9.56 19.63 40.59 9.50 19.60 86.79
-Strengths 42.93 11.13 22.09 43.36 11.20 22.44 87.84
-Summary 45.65 13.60 25.85 46.67 13.92 26.45 88.23
-Mistral $(5,5)$ Questions 39.20 9.00 19.91 40.67 9.32 20.83 86.09 6.44
-Weaknesses 40.27 9.21 19.46 39.49 9.01 19.13 86.71
-Strengths 43.45 11.49 22.34 44.11 11.61 22.82 87.84
-Summary 46.24 14.44 26.32 46.56 14.57 26.48 88.33
-Llama $(5,5)$ Questions 39.54 9.31 20.01 38.43 9.02 19.62 86.25 5.30
-Weaknesses 40.72 9.66 19.79 40.72 9.59 19.93 86.77
-Strengths 43.43 11.76 22.24 43.36 11.69 22.34 87.91
-Summary 46.92 14.90 26.74 47.30 14.99 26.92 88.36
-Table 5: The sectionwise performance across the ICLR-24 reviews test dataset after finetuning Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral).
-Per section of the review, the best results have been made bold, while the second-best results have been underlined.
-Please note that the GPU configurations for finetuning Llama and Mistral were different.
-For details, refer to Appendix C.
-As shown in Table 5, our best-performing configuration, Llama $(5,5)$ consistently outperforms the baselines across ROUGE-F1, ROUGE-Recall, and BERTScore metrics.
-Llama $(5,5)$ configuration is finetuned with input passages inferred from the GNN $(5,5)$.
-This highlights its ability to generate reviews that closely align with gold-standard references.
+To ensure that these gains are not solely due to using a more capable LLM, we also evaluate our method using Mistral.
+The results demonstrate that even with Mistral, integrating our Multi Modal GNN-based passage selection method leads to substantial improvements, confirming that the performance boost is primarily due to our graph-based retrieval strategy rather than the choice of the LLM alone.
 The integration of a Graph Attention Network (GAT) enables the model to effectively capture complex dependencies between relevant passages and sentences within the paper, offering the essential context required to generate comprehensive reviews.
-Notably, even the fastest variant of our fine-tuned models, Llama $(5,3)$, which requires the shortest training time, performs competitively.
-Despite using a smaller number of retrieved passages, this variant shows only an average marginal decrease of 0.71% in performance metrics compared to the best-performing configuration, Llama $(5,5)$.
-Further, it consistently outperforms the state-of-the-art (SOTA) baselines, while reducing training time by 32.07%.
-This trend is consistently observed across all metrics for Mistral too.
-It highlights that our GNN-based approach can achieve near-optimal performance with significantly less computational overhead.
-As shown in Table 5, a detailed analysis of our model’s performance across different review sections reveals that the Summary and Strengths sections achieve higher scores across ROUGE and BERTScore in our best-performing model, Llama $(5,5)$.
-This indicates that the model is effective at capturing and utilizing essential context from the paper to generate these key components of the review.
-Conversely, the Weaknesses and Questions sections receive comparatively lower scores.
-This indicates that identifying limitations and raising questions require a deeper semantic understanding of the paper as weaknesses may not always be explicitly stated in the passages of the paper.
-Overall, GNN helps to reflect on the content of the paper comprehensively to generate suitable reviews.
-5.3 LLM-as-a-judge Evaluation
-Review Review Quality Assessment Criteria
-Reviewer Confidence Review Thoroughness Review Constructiveness Review Helpfulness
+Our fastest, second-best, fine-tuned model, Llama $(5,3)$, performs competitively while being highly efficient.
+It reduces the training time by 32.07% with only a 0.71% drop in performance compared to our best model, consistently outperforming all baselines.
+This shows our GNN-based approach achieves near-optimal performance with significantly less computational overhead.
+5.2. Aspect Coverage Analysis
+As discussed by Yuan et al.
+(2022), each comment in a peer review typically addresses diverse aspects of an academic research paper, such as Summary (SUM), Motivation/Impact (MOT), Originality (ORI), Soundness/Correctness (SOU), Substance (SUB), Replicability (REP), Meaningful Comparison (CMP), and Clarity (CLA).
+These cannot be reliably assessed by conventional automatic metrics like ROUGE-F1 or BERTScore.
+To address this, Yuan et al.
+(2022) introduces an Aspect Tagger which can be used to compute Aspect Coverage (ACov) that measures the proportion of aspects covered by a peer-review.
+We evaluate the two best-performing models Llama $(5,5)$ and Llama $(5,3)$ using this tagger.
+Table 5.
+Aspect Coverage (%) of model-generated reviews evaluated using the Aspect Tagger (Yuan et al., 2022).
+Summary (SUM), Clarity (CLA), Soundness/Correctness (SOU), Substance (SUB), Originality (ORI), Meaningful Comparison (CMP), Replicability (REP), and Motivation/Impact (MOT).
+Model SUM CLA SOU SUB ORI CMP REP MOT
+Llama (5,3) 100.0 99.6 98.9 95.2 92.6 79.0 72.0 66.8
+Llama (5,5) 100.0 99.3 99.3 95.9 90.0 84.5 71.6 66.8
+As shown in Table 5, both variants achieve strong coverage for Summary, Clarity, and Soundness, indicating the reliability of our model-generated reviews in capturing these core aspects.
+However, the aspect coverage for Motivation/Impact and Replicability is comparatively lower.
+Overall, the results suggest that our model captures most key aspects, indicating that the generated reviews are comprehensive and address the quality of multiple dimensions of a research paper.
+5.3. Human Evaluation
+Understanding the limitations of automatic evaluation metrics as highlighted by Hanna and Bojar (2021) and Schluter (2017), we conduct human evaluations to assess the acceptability of the generated reviews.
+We model our evaluation around two distinct aspects of review quality.
+The first is a Qualitative Assessment, where the generated review is evaluated in isolation across four key criteria: Reviewer Confidence, Review Thoroughness, Review Constructiveness, and Review Helpfulness.
+The second is a Factual Validity Assessment, which ensures the review’s specific claims are factually grounded in the source document.
+Since human evaluation is very costly, we conduct each of the two tasks on a subset of the test dataset, as described in their individual subsections.
+Additionally, we only considered Llama $(5,3)$ and Llama $(5,5)$, our best performing models for the evaluation process.
+5.3.1. Qualitative Assessment
+A total of 30 papers are chosen at random from our test set.
+Each of these papers are associated with 3 reviews, one Gold, one from Llama $(5,3)$, and one from Llama $(5,5)$, thus, resulting in a total set of 90 reviews.
+A total of 27 annotators are chosen, and each annotator is given 10 reviews to annotate.
+The evaluation criteria is designed to capture key aspects of review quality and include:
+  1. (1)
+    Reviewer Confidence (RCf): Reflects the reviews’ depth of technical understanding and engagement with the content.
+  2. (2)
+    Review Thoroughness (RT): Measures how well the review covers the methodology, results, and reproducibility.
+  3. (3)
+    Review Constructiveness (RCt): Assesses the clarity and usefulness of suggestions for improvement.
+  4. (4)
+    Review Helpfulness (RH): Indicates how well the review identifies important issues and explains them clearly to the authors.
+Each annotator is tasked to assign a score from 1 to 5 for each of the above criteria.
+The evaluation guidelines are detailed in Appendix D.
+We end up with each review being annotated by 3 annotators.
+Score for a review is averaged across all its annotators, and the final score of a model, is the average score across all its reviews.
+The results of this experiment are summarized in Table 6.
+Table 6.
+Results of our human qualitative assessment.
+This table shows average scores (on a scale of 1-5) comparing gold-standard reviews with those generated by our two best-performing Llama models across four key criteria: Reviewer Confidence (RCf), Review Thoroughness (RT), Review Constructiveness (RCt), and Review Helpfulness (RH).
+Review RCf RT RCt RH
+Gold 3.54 3.49 3.26 3.41
+Llama (5,3) 3.37 3.30 3.27 3.34
+Llama (5,5) 3.27 3.24 3.19 3.30
+5.3.2. Factual Validity Assessment
+A total of 10 papers are chosen at random from our test set.
+Each of these papers are associated with 3 reviews, one Gold, and two generated by our best performing models.
+Thus, the final evaluation set consists of 10 papers, each associated with 3 reviews, one Gold, one from Llama $(5,3)$, and one from Llama $(5,5)$.
+These 10 papers are circulated among 10 annotators.
+It is ensured that a single paper is annotated by at least 2 annotators.
+Annotators are instructed to evaluate each statement within a review based on its factual grounding in the full source paper.
+For each point, they assign a rating from a four-point scale: ‘Completely Agree’, ‘Mostly Agree’, ‘Mostly Disagree’, or ‘Completely Disagree’.
+The evaluation guidelines are detailed in Appendix D.
+The results of this experiment are summarized in Table 7 and Table 8.
+Table 7.
+Detailed results of the Factual Validity Assessment.
+This table presents the percentage distribution of human agreement ratings for claims made in reviews from our Llama models and the gold standard.
+The abbreviations are: Completely Agree (CA), Mostly Agree (MA), Mostly Disagree (MD), and Completely Disagree (CD).
+Review CA MA MD CD
+Gold 39.12% 37.85% 16.40% 6.62%
+Llama (5,3) 32.64% 31.94% 22.92% 12.50%
+Llama (5,5) 32.27% 32.98% 23.05% 11.70%
+Table 8.
+Detailed results of the Factual Validity Assessment.
+This table shows the overall Positive Rating Percentage, calculated by summing the “Completely Agree” and “Mostly Agree” categories from Table 7, which indicates the factual soundness of the reviews.
+Review Positive Rating %
+Gold 76.97%
+Llama (5,3) 65.25%
+Llama (5,5) 64.58%
+5.4. LLM-as-a-Judge Evaluation
+As mentioned by Zheng et al.
+(2023), the LLM-as-a-Judge method offers an explainable and affordable way to approximate human preferences, which are otherwise costly to gather.
+We employ a large language model, Gemini-2.0-Flash [^9] https://deepmind.google/technologies/gemini/ as an automated judge to evaluate the quality of reviews generated by our proposed model as compared to the gold-standard reviews.
+For the LLM-as-a-judge task, the evaluation criteria and prompting guidelines are adapted from those used in our human qualitative assessment, which are detailed in Section 5.3.1.
+Since LLM-as-a-Judge evaluation is scalable and much cheaper to undertake, we evaluate all our models against the gold review, over the entire test dataset.
+A key finding from our evaluation is the strong correlation between the LLM-as-a-Judge scores (Table 9) and the results from our human annotators (Table 6).
+This alignment suggests that the LLM-as-a-Judge is a reliable and valid proxy for assessing review quality in this domain.
+Table 9.
+Qualitative evaluation of model-generated reviews using Gemini-2.0-Flash as the LLM judge.
+The table presents the average scores (out of 5) for all model configurations and the gold standard across four criteria: Reviewer Confidence (RCf), Review Thoroughness (RT), Review Constructiveness (RCt), and Review Helpfulness (RH).
+Bold values indicate the best score in each column (excluding Gold), and underlined values indicate the second-best.
+Review RCf RT RCt RH
 Gold 3.98 3.87 3.79 3.79
@@
 Llama $(5,5)$ 3.63 3.44 3.47 3.41
-Table 6: The table reports the qualitative evaluation of model-generated reviews using Gemini-2.0-Flash as an LLM judge.
-It presents the average scores (out of 5) across four review quality criteria: Reviewer Confidence, Review Thoroughness, Review Constructiveness, and Review Helpfulness, in comparison to gold reviews on the test set.
-Bold highlights the model best scores, and underlined values indicate the second-best.
-As mentioned by Zheng et al.
-(2023), the LLM-as-a-judge method offers an explainable and affordable way to approximate human preferences, which are otherwise costly to gather.
-We employ a large language model, Gemini-2.0-Flash [^4] https://deepmind.google/technologies/gemini/ as an automated judge to evaluate the quality of reviews generated by our proposed model as compared to the gold-standard reviews.
-The evaluation criteria are designed to capture key aspects of review quality and include:
-  1. 1.
-    Reviewer Confidence: reflects the reviewer’s depth of technical understanding and engagement with the content.
-  2. 2.
-    Review Thoroughness: measures how well the review covers the methodology, results, and reproducibility.
-  3. 3.
-    Review Constructiveness: assesses the clarity and usefulness of suggestions for improvement.
-  4. 4.
-    Review Helpfulness: indicates how well the review identifies important issues and explains them clearly for the authors.
-The model is prompted to assign a score from 1 to 5 based on the above criteria.
-The evaluation prompt is detailed in Appendix D.
-As shown in Table 6, our best-performing model, Llama $(5,5)$, scores only 9.59% below gold reviews across all criteria.
-The fastest variant, Llama $(5,3)$, achieves comparable quality with just a 9.78% drop relative to gold reviews while reducing training time by 32.07%.
-All Llama and Mistral variants consistently score above 3.0, indicating that the reviews generated grasp key contributions and provide relevant feedback with actionable suggestions.
-5.4 Ablation Studies
+5.5. Ablation Studies
 We perform ablation studies by systematically varying the values of $k\in\{1,3,5\}$ and $m\in\{3,5\}$ during LLM fine-tuning.
@@
 Additionally, we observe that higher values of $m$ result in longer training times, as the GNN retrieves and passes a larger set of passages as input for finetuning the LLMs.
-6 Conclusion
-In this work, we introduce AutoRev, an automatic peer review generation system designed specifically for academic research papers.
-Our approach leverages a novel graph-based framework to represent the structure of a paper which enables the effective extraction of important passages using a Graph Neural Network.
-This strategy of selecting passages and reducing the input length to LLMs addresses the limitations of traditional methods, like finetuning using the entire paper as input Yu et al.
-(2024), which leads to sub-optimal performance of LLMs due to large input token length Hsieh et al.
-(2024).
-Unlike DPR-based methods that require initial grounding on human written references to retrieve relevant passages, AutoRev employs a GNN that learns to identify important content directly from the paper’s graphical representation, without the need for grounding on existing reviews.
-This not only streamlines the process but also reduces computational overhead while enhancing review quality.
-Our experiments on the ICLR 2024 dataset demonstrate that AutoRev consistently outperforms existing baselines, by 58.72% on popular evaluation metrics like ROUGE, and BERTScore.
-Overall, AutoRev provides a promising solution for pre-review feedback, aiding both authors and reviewers in the academic publication process.
-Future work will explore extending this framework to a broader range of academic domains and investigating further improvements in graph construction and neural retrieval techniques to continue advancing the state-of-the-art in automated review generation.
-Limitations
-While our framework provides a robust and high-performing solution for review generation, we acknowledge that no automatic system can fully replicate human judgment in complex tasks such as peer review, as noted by Ye et al.
+6. Limitations
+While our framework provides a robust and high-performing solution for review generation, we acknowledge that no automatic system can fully replicate human judgment in complex tasks such as peer reviewing, as noted by Ye et al.
 (2024).
 Our goal is to assist authors in refining their papers and we aim for our system to support human reviewers.
-Currently, our system is primarily focused on ICLR papers due to the limited availability of public datasets.
+Currently, our system is primarily focused on ICLR 2024 papers due to the limited availability of public datasets containing consolidated reviews.
 However, we believe in the generalization capabilities of our model, given the proven effectiveness of graph-based approaches in various downstream NLP tasks, as discussed in Section 1.
-Ethical Considerations
+Our reliance on human written reviews, consolidated using GPT-4 represents another limitation.
+While multi-review datasets more closely reflect the diversity of real peer review, they are often noisy and contradictory, which can hinder model learning.
+Consolidated reviews provide a clearer and more consistent supervision signal, but at the cost of reducing the richness of differing reviewer perspectives.
+Future work should explore hybrid approaches that retain both consistency and diversity.
+7. Ethical Considerations
 The deployment of AutoRev raises key ethical concerns related to bias, misuse, and academic integrity.
@@
 To address these issues, we advocate for responsible usage of AutoRev as a supplementary tool, accompanied by transparency in AI involvement and safeguards to preserve academic integrity.
+8. Conclusion
+In this work, we introduce AutoRev, an automatic peer review generation system designed specifically for academic research papers.
+Our approach leverages a novel graph-based framework to represent the structure of a paper which enables the effective extraction of important passages using a Graph Neural Network.
+This strategy of selecting passages and reducing the input length to LLMs addresses the limitations of traditional methods, like finetuning using the entire paper as input (Yu et al., 2024), which leads to sub-optimal performance of LLMs due to large input token length (Hsieh et al., 2024).
+Unlike DPR-based methods that require initial grounding on human written references to retrieve relevant passages, AutoRev employs a GNN that learns to identify important content directly from the paper’s graphical representation, without the need for grounding on existing reviews.
+This not only streamlines the process but also reduces computational overhead while enhancing review quality.
+Our experiments on the ICLR 2024 dataset demonstrate that AutoRev consistently outperforms existing baselines, by 58.72% on popular evaluation metrics like ROUGE, and BERTScore.
+These findings are further strongly backed by human evaluations, where the generated reviews were rated nearly as constructive and acceptable as human-written ones, suggesting AutoRev could meaningfully reduce reviewer workload.
+Beyond these performance gains, we emphasize AutoRev’s broader role in the web ecosystem.
+Academic research papers represent some of the most authoritative and trusted content on the web.
+By accelerating the peer-review workflow and enabling scalable, automated feedback, AutoRev contributes to the timely dissemination of credible scholarly knowledge online.
+In doing so, it supports not only authors and reviewers but also the wider web community that depends on accurate, trustworthy research for discovery, indexing, and downstream applications.
+Future work will explore extending this framework to a broader range of academic domains and investigating further improvements in graph construction and neural retrieval techniques to continue advancing the state-of-the-art in automated review generation.
 Appendix A Ablation studies
@@
 While calculating the ROUGE metrics, we enable the use_stemmer option to ensure consistent matching of words with different morphological forms, to improve the reliability of the evaluation.
+Table 10.
+Overall performance of Mistral across different $(k,m)$ settings.
+Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours).
+Bold highlights the best scores, and underlined values indicate the second-best.
 Model ROUGE-F1 ROUGE-Recall BERTScore F1 Training Time (hrs)
@@
 Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
-Table 7: Overall performance of Mistral across different $(k,m)$ settings.
+Table 11.
+Overall performance of Llama across different $(k,m)$ settings.
 Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours).
@@
 Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
-Table 8: Overall performance of Llama across different $(k,m)$ settings.
-Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours).
-Bold highlights the best scores, and underlined values indicate the second-best.
 Appendix B GNN Implementation Details
-We trained and tested our model on a single NVIDIA RTX6000 GPU with 48GB of VRAM.
-B.1 Model Architecture
+We train and test our model on a single NVIDIA RTX6000 GPU with 48GB of VRAM.
+B.1. Model Architecture
 The model contains a Graph Neural Network (GNN) module and a passage classifier.
@@
 This is followed by an ELU activation and dropout ($p=0.2$).
-The second GATConv layer maps the $128\times 4$ input to 256-dimensional features using 2 attention heads (output dimension $256\times 2$).
+The second GATConv layer maps the $128\times 4$ input to 256 dimensional features using 2 attention heads (output dimension $256\times 2$).
 A residual connection similarly projects the prior layer’s output to $256\times 2$, followed by ELU and dropout ($p=0.2$).
@@
 The classifier is implemented as a multi-layer perceptron (MLP) comprising the following sequence of operations:
-  • A linear transformation projecting the input to a 512-dimensional space.
+  • A linear transformation projecting the input to a 512 dimensional space.
   • A SiLU activation followed by dropout with a probability of $p=0.3$.
@@
 This architecture enables the model to learn nonlinear transformations of the graph-encoded passage representations while incorporating regularization and normalization for stable training.
-B.2 Training
+B.2. Training
 The model is trained for 30 epochs, with the best validation performance checkpoint retained for inference. The loss function used for training is the Cross Entropy Loss.
@@
 The full set of hyperparameters used during training is listed in Table 12.
+Table 12. Hyperparameters Used During Finetuning
 Hyperparameter Value
@@
 LR Scheduler Type cosine
-Table 9: Hyperparameters Used During Finetuning
-Prompt for Finetuning
-The prompt used for finetuning both Llama 3.1 8B and Mistral-Instruct-v0.2 can be found below:
-Prompt Used During Finetuning Below is an instruction that describes a task, paired with an input that provides further context.
-Write a response that appropriately completes the request.
-### Instruction:
-Generate a structured review for the research
-paper passages provided below.
-The review should include a summary of the paper, its strengths, weaknesses, and questions for the authors.
-Consider that the paper is being reviewed for the ICLR conference.
-### Research Paper Passages:
-### Review for the paper:
+C.1. Prompt for Finetuning
+The prompt used for fine-tuning both Llama 3.1 8B and Mistral-Instruct-v0.2 is shown in Listing C.1.
 […]
-[…]
-[…]
-[…]
-The extracted relevant passages are inserted ahead of the "Research Paper Passages" heading.
+The extracted relevant passages are inserted ahead of the ”Research Paper Passages” heading.
 Similarly, the corresponding content for each review section (Summary, Strengths, Weaknesses, and Questions) is added after their respective headings in the generated review.
-Appendix D LLM-as-a-judge evaluation
-** Instruction ** : You are an expert meta-reviewer. Evaluate the following peer review text according to the scoring
-[…]
-Give rating for the categories described below on the full scale from 1(lowest rating) to 5(highest rating).
-Assign scores independently per category (they may differ significantly).
-**Category 1: Reviewer Confidence**
-1 – Unqualified: Admits lack of expertise | Fundamental misunderstandings | Fails to engage with technical content
-2 – Limited: Surface-level comprehension | Misses >2 key concepts | Confuses established methods
-3 – Functional: Grasps main contributions | Identifies 1-2 technical flaws | Limited literature context
-4 – Qualified: Precise technical critique | Contextualizes in recent work (>3 citations) |
-Identifies methodological tradeoffs
-5 – Authoritative: Demonstrates field leadership | Predicts follow-up research directions |
-Reveals fundamental limitations in approach
-**Category 2: Review Thoroughness**
-1 – Negligent: <3 substantive comments | Skips methodology/results | Fails to address claims
-2 – Patchy: Only addresses obvious aspects | 0 analysis of experiments | No ethical consideration
-3 – Standard: Checks main claims | Basic methodology review | 1 reproducibility concern noted
-4 – Rigorous: Line-by-line methodology check | Statistical validity analysis | 2+ improvement suggestions
-5 – Forensic: Alternative approach proposals | Re-analysis of key results |
-Compliance checklist (ethics, reproducibility, etc.)
-**Category 3: Constructiveness**
-1 – Destructive: Personal attacks | Unsubstantiated dismissal | 0 actionable feedback
-2 – Demanding: Unrealistic requests ("redo all experiments") | Vague quality complaints
-3 – Directive: General improvement areas | 1-2 specific examples | Mixed positive/negative tone
-4 – Strategic: Priority-ranked suggestions | Alternative methodology paths | Conference-specific improvements
-5 – Transformative: Step-by-step revision roadmap | Code/experiment snippets provided | Before-after examples
-**Category 4: Helpfulness**
-1 – Misleading: Factually incorrect suggestions | Contradicts paper’s goals | Harmful recommendations
-2 – Abstract: "Improve writing" without examples | "Add experiments" without design
-3 – Guided: Section-specific edits | 2-3 implementable suggestions | Partial error diagnosis
-4 – Impactful: Directly address weaknesses | Templates/examples provided | Clear priority levels
-5 – Revelatory: Exposes fundamental flaws | Provides benchmark comparisons | Supplies missing literature
-**Scoring Methodology**
-For EACH category:
-1. Find the HIGHEST level (5-1) where ALL criteria are FULLY satisfied (Base Level)
-2. Check criteria in the NEXT HIGHER LEVEL (Base Level + 1):
-   - Calculate partial credit: (Satisfied Criteria in Higher Level / Total Criteria in Higher Level)
-3. Final Score = Base Level + Partial Credit
-4. Round to 2 decimal places (0.33 increments for 3-criteria levels, 0.5 for 2-criteria levels)
-[…]
-[…]
-Respond ONLY with this JSON format:
-  "review_scores": {{
-    "Category 1": "",
-    "Category 2": "",
-    "Category 3": "",
-    "Category 4": ""
-Figure 4: Prompt used during evaluating the quality of generated reviews with LLM as a Judge
-The prompt used for LLM-as-a-judge with Gemini-2.0-Flash is shown in Figure 4. The corresponding model-generated review is inserted under the appropriate subheadings within the prompt.
-To demonstrate the effectiveness and sensitivity of our evaluation prompt, we include two manually crafted sample reviews: one representing a poorly written, low-quality review, and another simulating an average review.
-These examples are not generated by our system but are intentionally designed to test whether the LLM-as-a-judge can distinguish between different qualities of feedback.
-This validation step supports the reliability of the evaluation framework in scoring model outputs.
-Example 1: Structured Critical Review (Moderate Quality)
-Summary: The paper tackles a topic that lacks clear novelty, presents a poorly motivated approach, and delivers incomplete experiments with unconvincing results.
-The writing is disorganized, and key methodological and comparative details are missing, making it difficult to assess the contribution.
-Strengths:
-  • Addresses a relevant area of graph-based modeling (though similar to existing work).
-  • Attempts to combine embeddings and graph structure in a single framework.
-Weaknesses:
-  • Motivation is vague and fails to establish why the problem matters.
-  • No thorough comparison to closely related methods; literature review is superficial.
-  • Methodology section is under-specified, lacking algorithmic details and hyperparameter settings.
-  • Experimental results are incomplete: missing error bars, statistical analysis, and ablation studies.
-  • Figures and tables are low resolution or poorly labeled, hindering interpretability.
-Questions:
-  • What is the core novelty compared to the most recent graph-embedding methods?
-  • How were key hyperparameters chosen, and how sensitive are the results to those choices?
-  • Can you provide ablation studies showing the contribution of each component in the model?
-  • Why are error bars and statistical significance tests omitted from the reported results?
-  • How does your method scale to larger datasets, and what is its computational complexity?
-The evaluation scores obtained for Example-1 are as follows,
-  1. 1.
-    Reviewer Confidence : 3.0
-  2. 2.
-    Review Thoroughness : 3.0
-  3. 3.
-    Review Constructiveness : 3.33
-  4. 4.
-    Review Helpfulness : 3.0
-Example 2: Informal and Unstructured Review (Low Quality)
-This paper is pretty weak.
-The writing is sloppy and hard to follow.
-I’m not convinced the idea is novel.
-There are lots of papers out there doing similar stuff.
-The experiments seem incomplete, and the results aren’t very impressive.
-I don’t think this will be of interest to ICLR readers.
-You should do more experiments or maybe scrap the whole thing and start over.
-Comments:
-  • The motivation is unclear.
-  • There’s no thorough comparison to related work.
-  • The methodology description is too brief.
-  • Results tables have no error bars.
-  • Figures are low resolution and hard to read.
-Recommendation: Reject.
-The evaluation scores obtained for Example-2 are as follows,
-  1. 1.
-    Reviewer Confidence : 2.0
-  2. 2.
-    Review Thoroughness: 1.0
-  3. 3.
-    Review Constructiveness : 2.0
-  4. 4.
-    Review Helpfulness : 2.0
+Appendix D Human Evaluation
+Due to space constraints, the detailed evaluation guidelines are made available in an anonymized GitHub repository: Anonymous Repository.
+For the camera-ready version, we will release these materials in a public GitHub repository to ensure full transparency and accessibility.
+We will now proceed to show the snapshots of our annotation UI, developed using Streamlit  [^10] https://streamlit.io/.
+D.1. Qualitative Assessment
+Examples of the user interface used for the Qualitative Assessment can be seen in Figure 4 and Figure 5.
+Figure 4.
+This interface component corresponds to the Qualitative Assessment task described in Section 5.3.1.
+It marks the beginning of the review presentation.
+Due to space constraints, only the summary and a subset of the strengths from one review are shown here.:
+Figure 5. This interface component corresponds to the task described in Section 5.3.1. In this part, annotators provide a final rating across the four categories and submit their annotation.:
+D.2. Factual Validity Assessment
+Examples of the user interface used for Factual Validity Assessment can be seen in Figure 6a and Figure 6b.
+(a) The original paper is displayed in a PDF viewer to provide factual context.
+(b) Each point of the review is rated for accuracy using a dropdown menu.
+Figure 6.
+The user interface for the Factual Validity Assessment task described in Section 5.3.2.
+On the left (a), the full research paper is presented to the annotator.
+On the right (b), the annotator rates each point of the review based on its consistency with the source document.:
+Appendix E Example of a Generated Review
+We present a review generated by our system for the paper: On the Role of Edge Dependency in Graph Generative Models [^11] https://openreview.net/forum?id=LCQ7YTzgRQ
+Summary
+The paper introduces a novel framework for evaluating generative models of graphs, focusing on the concept of “overlap” to ensure both accuracy and edge diversity.
+It categorizes graph generative models into three levels of complexity: edge independent, node independent, and fully dependent models.
+The paper derives theoretical bounds on triangle counts and other short-length cycles based on overlap and provides new generative models for each level.
+However, the paper’s novelty and practical contributions are questioned due to the use of existing methods and the lack of significant empirical validation.
+The experimental section is seen as limited and does not convincingly demonstrate the superiority of the proposed methods over existing ones.
+Strengths
+  • The paper introduces a novel evaluation framework that considers the importance of model-generated graph overlap to ensure both accuracy and edge diversity, which is a significant contribution to the field of graph generative models.
+  • The paper presents a hierarchy of graph generative models categorized into three levels of complexity, which encapsulates a wide range of prevalent methods and provides a structured framework for evaluating graph generative models.
+  • The theoretical bounds on the number of triangles and other short-length cycles are well-presented and support the findings with instances demonstrating the asymptotic optimality of these bounds.
+  • The authors introduce new generative models for each of the three hierarchical levels, leveraging dense subgraph discovery (Gionis & Tsourakakis, 2015), which is a practical application of the theoretical findings.
+Weaknesses
+  • The novelty of the paper is limited as it primarily relies on existing methods and theoretical frameworks, and the experimental results do not convincingly demonstrate the superiority of the proposed methods over existing ones.
+  • The empirical evaluation is weak and does not provide sufficient evidence to support the claims made in the paper.
+The experimental results are not robust and the paper does not provide enough comparisons with other models to establish the effectiveness of the proposed methods.
+  • The paper has several presentation issues, including unclear definitions and notations, which could hinder the understanding of the content for readers not familiar with graph generative models.
+  • The practical applications and the significance of the proposed methods are not clearly demonstrated, and the paper lacks a detailed discussion on the limitations and potential future work.
+Questions
+  • Can the authors clarify the definitions and notations used in the paper, particularly the ones related to the overlap and triangle density, to ensure a clearer understanding of the content?
+  • How do the authors justify the novelty of their work given the reliance on existing methods and theoretical frameworks?
+  • What specific contributions does the paper make to the field of graph generative models, and can these be clearly articulated?
+  • Why was the empirical evaluation conducted in a limited manner, and how can the paper be improved to provide more robust and convincing experimental results?
+  • Could the authors discuss the practical applications of the proposed methods and the significance of their work in more detail?
+  • Are there any plans to extend the theoretical bounds to higher-order cycles, and how might this affect the overall conclusions drawn in the paper?

### v3

@@
-AutoRev: Multi-Modal Graph Retrieval for Automated Peer-Review Generation
-Abstract.
-Enhancing the quality and efficiency of academic publishing is critical for both authors and reviewers, as research papers are central to scholarly communication and a major source of high-quality content on the web.
-To support this goal, we propose AutoRev, an automatic peer-review system designed to provide actionable, high-quality feedback to both reviewers and authors.
-AutoRev leverages a novel Multi-Modal Retrieval-Augmented Generation (RAG) framework that combines textual and graphical representations of academic papers.
-By modelling documents as graphs, AutoRev effectively retrieves the most pertinent information, significantly reducing the input context length for LLMs and thereby enhancing their review generation capabilities.
-Experimental results show that AutoRev outperforms state-of-the-art baselines by up to 58.72% and demonstrates competitive performance in human evaluations against ground truth reviews.
-We envision AutoRev as a powerful tool to streamline the peer-review workflow, alleviating challenges and enabling scalable, high-quality scholarly publishing.
-By guiding both authors and reviewers, AutoRev has the potential to accelerate the dissemination of quality research on the web at a larger scale.
+Graph-Guided Passage Retrieval for Author-Centric Structured Feedback
+Abstract
+Obtaining high-quality, pre-submission feedback is a critical bottleneck in the academic publication lifecycle for researchers.
+We introduce AutoRev, an automated author-centric feedback system that generates structured, actionable guidance prior to formal peer review.
+AutoRev employs a graph-based retrieval-augmented generation framework that models each paper as a hierarchical document graph, integrating textual and structural representations to retrieve salient content efficiently.
+By leveraging graph-based passage retrieval, AutoRev substantially reduces LLM input context length, leading to higher-quality feedback generation.
+Experimental results demonstrate that AutoRev significantly outperforms baselines across multiple automatic evaluation metrics, while achieving strong performance in human evaluations.
 Code will be released upon acceptance.
-Graph Neural Networks (GNNs), Retrieval Augmented Generation (RAG), Large Language Models (LLMs), Automatic Peer Reviewing
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Information extractionccs: Computing methodologies Natural language generationccs: Information systems Language modelsccs: Information systems Document structure
-1. Introduction
-Figure 1. Example document graph with nodes for headings, sub-headings, passages, and sentences. This structure allows AutoRev to capture hierarchical and sequential dependencies across sections.:
-Recent research has increasingly focused on developing automatic systems for peer review, including tools that assist reviewers, assess peer-review processes, and directly generate reviews for academic papers (Yu et al., 2024; Aitymbetov and Zorbas, 2025; D’Arcy et al., 2024; Idahl and Ahmadi, 2025).
-Concurrently, there has been growing interest in examining the strengths and limitations of such systems, particularly the use of large language models (LLMs) for automated review generation (Ye et al., 2024).
-Motivated by both of these trends, we propose AutoRev, a system designed to generate peer reviews from a guidance perspective, offering constructive feedback to support researchers and reviewers.
-With the rise of Graph Neural Networks, the use of graphs as extractors or as guides for facilitating downstream tasks has recently gained traction in the NLP community (Chitale et al., 2025; Gao et al., 2025; Agarwal et al., 2025; Zhu et al., 2025).
-Inspired by this, we propose to use graphs in the domain of academic paper analysis, specifically, in generating reviews that closely replicate the structure and quality of those produced by existing peer review systems in conferences.
-Furthermore, by leveraging the graph’s capability as an extractor of passages, we significantly reduce the input token length to the LLM.
-Reducing the token length of the input is shown to enhance LLM performance (Hsieh et al., 2024).
+Graph-Guided Passage Retrieval for Author-Centric Structured Feedback
+Maitreya Prafulla Chitale1thanks: Equal contribution.
+Ketaki Mangesh Shetye [^1] footnotemark: 1 Harshit Gupta2Manav Chaudhary3Manish Shrivastava1Vasudeva Varma1 1IIIT Hyderabad  2Microsoft, India  3Sentisum {maitreya.chitale, ketaki.shetye}@research.iiit.ac.in guptaharshi@microsoft.com manav@sentisum.com {m.shrivastava, vv}@iiit.ac.in
+1 Introduction
+Recent advancements have focused on assisting reviewers and directly generating peer reviews Yu et al.
+(2024); Aitymbetov and Zorbas (2025); D’Arcy et al.
+(2024); Idahl and Ahmadi (2025).
+Concurrently, there has been growing interest in examining the strengths and limitations of such systems, particularly the use of large language models (LLMs) in this domain Ye et al.
+(2024).
+Relying on LLMs to generate or inform final review decisions remains risky due to issues of bias, overconfidence, and superficial feedback and these decision-oriented tools lack utility for authors in the pre-submission phase.
+Consequently, we advocate for a new paradigm towards author-centric support.
+We introduce AutoRev, an author-centric system that generates structured, actionable pre-submission feedback containing Summary, Strengths, Weaknesses, and Questions, without producing acceptance recommendations or replacing human peer review.
+Figure 1: Example document graph with nodes for headings, sub-headings, passages, and sentences. This structure allows AutoRev to capture hierarchical and sequential dependencies across sections.
+With the rise of Graph Neural Networks, the use of graphs as extractors or as guides for facilitating downstream tasks has recently gained traction in the NLP community Chitale et al.
+(2025); Gao et al.
+(2025); Agarwal et al.
+(2025); Zhu et al.
+(2025).
+Inspired by this, we leverage the graph’s capability as an extractor of passages to significantly reduce the input token length to the LLM.
+Reducing the token length of the input is shown to enhance LLM performance Hsieh et al.
+(2024).
 To summarize our main contributions:
-  1. (1)
-    We propose a system for generating concise, structured reviews (Summary, Strengths, Weaknesses, and Questions) to assist authors and reviewers.
-To our knowledge, this is the first approach which uses graphs in this domain.
-  2. (2)
-    Our approach improves over SoTA baselines by 58.72% on average across all metrics while being computationally efficient.
-  3. (3)
-    We conduct thorough human evaluations and other automatic evaluations to demonstrate the competitive performance of our method against ground truth reviews.
-It is important to highlight that this paper does not seek to determine the final reviews of a paper.
-Rather, we present our framework as a tool to provide timely feedback to researchers and serve as a valuable guide for reviewers.
-We envision AutoRev as a step towards scalable, automated peer-review systems that can complement human judgment.
-With the continued surge in conference submissions across disciplines, the demand for such tools is increasingly pressing [^1] https://aaai.org/aaai-launches-ai-powered-peer-review-assessment-system/.
-By alleviating the bottlenecks of the traditional peer-review process, AutoRev can help accelerate the dissemination of high-quality research on the web, ensuring that credible scholarly knowledge remains accessible at scale.
-2. Related Works
-A significant amount of research has been conducted on aiding peer reviews for academic papers.
-Some studies explore the involvement of AI in the reviewer selection process (Aitymbetov and Zorbas, 2025), while others focus on generating meta-reviews (Kumar et al., 2023).
-Prior work has also examined how AI/NLP techniques can contribute to various aspects of the peer-review pipeline (Kuznetsov et al., 2024).
-Faizullah et al.
-(2024) explore a DPR-based approach for generating suggestive limitations, highlighting the necessity of retrieval methods to overcome length constraints.
-Building on this, we use DPR outputs as the training data for our Graph Neural Network (GNN), enabling it to learn to identify salient passages for review generation.
+  1. 1.
+    We propose an author-centric system for generating concise, structured pre-submission feedback for academic papers.
+  2. 2.
+    We introduce a graph-based retrieval framework that learns to identify salient passages, reducing LLM input length while improving feedback quality.
+  3. 3.
+    We validate our approach using comprehensive human evaluations to assess feedback quality and factual grounding, in addition to standard automatic metrics.
+2 Leveraging Peer Reviews as Proxy for Pre-Submission Feedback
+In the absence of large-scale, publicly available datasets explicitly designed for pre-submission feedback, we use human-written peer reviews as a proxy, as they provide expert critiques of writing, methodology, and presentation aligned with the guidance authors seek prior to submission.
+In particular, the Weaknesses and Questions sections offer actionable critiques and clarifications.
+Since the dataset spans papers at varying stages of writing maturity, the associated reviews reflect feedback targeted at authors with differing levels of experience.
+We further support this choice through an impact-oriented user study (Section 6.3.4).
+3 Related Works
+Recent work has explored the use of NLP techniques to support scholarly writing and review workflows, including reviewer selection Aitymbetov and Zorbas (2025), meta-review generation Kumar et al.
+(2023), and automated review generation Yu et al.
+(2024).
+Many of these approaches rely on large language models, ranging from modular agent-based systems D’Arcy et al.
+(2024) to fine-tuned review generators Idahl and Ahmadi (2025).
+In parallel, a separate line of work focuses on author-assistive tools for specific writing sub-tasks, such as limitation suggestion Faizullah et al.
+(2024), citation recommendation and literature discovery Gu and Hahnloser (2023), and drafting individual sections like related work Hu and Wan (2014) or introductions Liebling et al.
+(2025).
+While these are effective within their scoped objectives, they are either reviewer-facing or narrowly targeted, and often operate on full-paper inputs at inference time.
+As a result, they raise concerns around hallucination, superficial feedback, computational cost, and overconfidence in model-generated judgments Ye et al.
+(2024), while failing to provide holistic, structured feedback that supports authors prior to submission.
 Yu et al.
-(2024) introduce an automated paper reviewing framework (SEA) that standardizes reviews using GPT-4222https://arxiv.org/abs/2303.08774 consolidation and proposes a self-correction strategy during fine-tuning.
-Ye et al.
-(2024) further caution that LLM-generated reviews may overrate papers or produce superficial feedback, reinforcing our design goal of providing a structured and constructive review rather than final accept/reject decisions.
-Among the most relevant recent systems, MARG (D’Arcy et al., 2024) employs a multi-agent framework, distributing different sections of a paper to specialized GPT-4 agents.
-This enables the model to handle long papers and produce more specific comments, but it comes at the cost of heavy computation and frequent inter-agent coordination errors.
-OpenReviewer (Idahl and Ahmadi, 2025) instead fully fine-tunes Llama-3.1-8B-Instruct on 79k reviews from ICLR and NeurIPS to generate structured reviews, adhering to the conference guidelines.
-This method of large-scale full fine-tuning is very costly.
-We adopt SEA (Yu et al., 2024) as our primary baseline since it is closest in scope to our work, representing a recent state-of-the-art framework for automated peer review generation on the same ICLR dataset.
-Other systems such as MARG and OpenReviewer differ substantially in design.
-MARG uses multi-agent GPT-4 orchestration, while OpenReviewer relies on large-scale fine-tuning, making direct benchmarking less meaningful.
-Thus, we benchmark against SEA for a fair and reproducible comparison.
-The integration of graph-based techniques into Retrieval Augmented Generation (RAG) (Lewis et al., 2020) is an emerging research area that aims to enhance the reasoning and factual grounding of LLMs.
-A recent comprehensive survey by Zhu et al.
-(2025) provides a foundational taxonomy for this field, categorizing systems based on their database construction, retrieval algorithms, and processing pipelines.
-Within this framework, AutoRev is a domain-specific Graph-RAG application that generates a hierarchical graph directly from a scientific paper’s text.
-It employs a learning-based Graph Attention Network (GAT) (Veličković et al., 2018) for retrieval that reduces the context length of input to LLMs, reducing the computational costs.
-3. Dataset
+(2024) introduce an automated paper reviewing framework (SEA) that standardizes reviews using GPT-4 OpenAI et al.
+(2024) consolidation and proposes a self-correction strategy during fine-tuning.
+We adopt SEA as our baseline as it is trained on the publicly available ICLR 2024 dataset and generates the same structured fields as AutoRev, among others.
+Studies have explored graph-based extensions to Retrieval-Augmented Generation (RAG) Lewis et al.
+(2020) to improve grounding and reasoning in LLMs.
+Zhu et al.
+(2025) provide a comprehensive taxonomy of Graph-RAG systems based on database construction, retrieval mechanisms, and processing pipelines.
+Within this framework, AutoRev represents a domain-specific Graph-RAG approach that constructs a hierarchical graph directly from a scientific paper and employs a learning-based Graph Attention Network veličković2018graph for passage retrieval.
+4 Dataset
 The dataset used in our study is borrowed from Yu et al.
 (2024), comprising papers submitted to various conferences along with multiple peer reviews per paper.
-Among these, the ICLR 2024 and NeurIPS 2023 datasets are publicly released, though the GPT-4–consolidated reviews for NeurIPS 2023 are not released.
+Among these, the ICLR 2024 and NeurIPS 2023 datasets are publicly released, though the GPT-4-consolidated reviews for NeurIPS 2023 are not available.
 Hence, our analysis focuses only on the ICLR 2024 papers.
-As reported in Yu et al.
-(2024), the dataset statistics for the ICLR 2024 papers are summarized in Table 1.
-Table 1. Dataset statistics for the ICLR 2024 dataset, as reported in Yu et al. (2024).
+Table 1 summarizes statistics for the ICLR 2024 papers Yu et al.
+(2024).
 Attribute Value
@@
 Domain Machine Learning
-Each paper has multiple peer reviews including sections such as Summary, Strengths, Weaknesses, Questions, Soundness, Presentation, Contribution, Confidence, and Rating to assess the quality and clarity of the paper.
-From the dataset, we utilize only the Summary, Strengths, Weaknesses, and Questions sections from each review, along with the full content of the corresponding paper.
-Other sections, such as Ratings are excluded, as our framework is designed to provide constructive feedback rather than produce a final decision.
-The selected sections offer qualitative assessments of the paper’s content and provide valuable insights to guide both authors and reviewers.
-For our finetuning process, we utilize the GPT-4 consolidated reviews dataset introduced by Yu et al.
-(2024).
-This dataset is constructed by using GPT-4 to consolidate multiple human-written peer reviews from OpenReview [^3] https://openreview.net/ into a single, unified review for each paper.
-4. Methodology
-Figure 2.
-The system begins with a parsed document from which DPR extracts passages for GNN training.
-The GNN learns to identify key passages relevant to review generation.
-The trained GNN selects important passages without requiring human-generated reviews.
+Table 1: Dataset statistics for the ICLR 2024 dataset, as reported in Yu et al. (2024).
+Each paper is associated with multiple reviews containing sections such as Summary, Strengths, Weaknesses, Questions, Soundness, Presentation, Contribution, Confidence, and Rating.
+We retain only the Summary, Strengths, Weaknesses, and Questions sections, as these encode constructive feedback for authors-centric pre-submission support (see Section 2), and exclude rating-based fields that correspond to final decision-making.
+For training, we use the GPT-4 consolidated reviews released by Yu et al.
+(2024), which merge multiple human-written reviews into a single structured review per paper.
+To assess cross-venue performance, we curate a supplementary dataset of 10 papers from COLM and NeurIPS 2025 (CNT_10).
+This set complements the ICLR 2024 test data (ICT) for the human evaluation detailed in Section 6.3.
+Figure 2: The system begins with a parsed document from which DPR extracts passages for GNN training.
+The GNN learns to identify key passages relevant to structured feedback generation.
+The trained GNN selects important passages without requiring human-generated feedback.
 These passages, combined with a structured prompt, are input to fine-tune an LLM.
-The fine-tuned LLM generates the final review efficiently, without external grounding.:
+The fine-tuned LLM generates the final structured feedback, without external grounding.
+5 Methodology
 In this section, we describe the implementation of our proposed framework.
@@
 The extracted passages are used as training data in Section 5.3, to train a Graph Neural Network (GNN) that learns to extract relevant passages from a research paper.
-The trained GNN enables our model to perform competitively on the test set without the need for grounding on human-generated reviews, as discussed in Section 4.4.
-Post that, we describe our process of fine-tuning LLMs to generate the final reviews, in Section 4.5.
+The trained GNN enables our model to perform competitively on the test set without the need for grounding on human-generated feedback, as discussed in Section 5.4.
+Post that, we describe our process of fine-tuning LLMs to generate the final structured feedback, in Section 5.5.
 An overview of our methodology can be found in Figure 2.
-4.1. DPR-based Passage Extraction
-Dense Passage Retrieval (DPR) is a neural retrieval model, primarily trained on question-answering datasets, that encodes both queries and context into dense vector representations.
-Given an input query, DPR retrieves the most relevant passages based on their semantic similarity in the dense embedding space.
-We use a pre-trained DPR model, to identify the most relevant passages corresponding to the different sections of the reviews that are Summary, Strengths, Weaknesses, and Questions, for each paper in the training set.
-Specifically, we use the context encoder [^4] https://huggingface.co/facebook/dpr-ctx_encoder-multiset-base and the query encoder [^5] https://huggingface.co/facebook/dpr-question_encoder-multiset-base from the multiset-base DPR variant.
+5.1 DPR-based Passage Extraction
+Dense Passage Retrieval (DPR) is a neural retrieval model trained primarily on question-answering datasets that encodes queries and passages into dense vector representations.
+Given an input query, DPR retrieves relevant passages based on semantic similarity in the embedding space.
+We use a pre-trained DPR model to identify passages corresponding to the Summary, Strengths, Weaknesses, and Questions sections of reviews for each paper in the training set.
+Specifically, we use the context encoder [^1] https://huggingface.co/facebook/dpr-ctx_encoder-multiset-base and query encoder [^2] https://huggingface.co/facebook/dpr-question_encoder-multiset-base from the multiset-base DPR variant.
 […]
-This final set of passages represents the most relevant excerpts from the full paper that are semantically linked to the review content.
+This final set of passages represents the most relevant excerpts from the full paper that are semantically linked to the feedback content.
 These extracted passages serve as the ground truth labels for training the GNN, as described in Section 5.3
-4.2. Graph Construction for Representing a Paper
+5.2 Graph Construction for Representing a Paper
 The idea of graph construction employed in our work is inspired by Zhao et al.
@@
 We borrow their idea of representing a document as a hierarchical graph but extend it beyond sentence level.
-We propose to build a deeper graph, which contributes towards learning more intricate inter-dependencies of passages and sentences in a paper.
-Specifically, we formalize the document structure as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where nodes $\mathcal{V}$ represent text segments with semantically encoded features, and $\mathcal{E}$ denote the hierarchical and sequential relationships that form the document’s structure.
+The graph is constructed directly from the parsed paper and is designed to generalize across diverse academic writing styles and section hierarchies.
+Formally, each paper is represented as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where nodes correspond to textual units at different granularities, including the paper root, section headings, sub-headings, passages, and sentences.
+Edges encode two complementary relationships: (i) hierarchical links ($\mathcal{E}_{hier}$) that reflect the document’s structural nesting and (ii) sequential links ($\mathcal{E}_{seq}$) that preserve the linear ordering of passages and sentences.
+Together, these connections allow information to propagate both vertically across document structure and laterally across adjacent text.
+A schematic illustration of the document graph is shown in Figure 1.
+Full details of node types, edge definitions, and construction procedures are provided in Appendix A.
+This design explicitly mitigates over-squashing in tree-like graphs, where exponential neighborhood growth leads to information bottlenecks at the root (Alon and Yahav, 2021; Topping et al., 2022).
+Following the rewiring strategies of Topping et al.
+(2022), we augment the hierarchical backbone with sequential edges to facilitate lateral information flow and reduce reliance on hierarchical connections.
+Additionally, the inherent structure of scientific papers results in graphs of shallow depth (typically $\leq 5$ hops), limiting long-range dependency compression and preserving fine-grained information from leaf nodes.
+5.3 GNN Training for Passage Retrieval
+To train a GNN on the graph constructed in Section 5.2, we take the data generated in Section 5.1, which maps each paper $p_{i}$ to a set of important passages, $\mathcal{O}_{i}=\{o_{1},o_{2},\dots,o_{n}\}$.
+Each passage is important for generating a feedback for the paper it is associated with.
+This mapping is used as a training set for training the GNN.
+The node embeddings for sentence nodes are initialized with embeddings from a Sentence-Transformer Reimers and Gurevych (2019) model, namely all-mpnet-base-v2.
+The node embeddings for all other nodes are initialized with zero embeddings of dimension $768$, matching that of the Sentence-Transformer model.
+We train a Graph Attention Network (GAT) veličković2018graph over the document graph to classify passage nodes as feedback-relevant or not.
+The GNN is trained using various configurations based on the values of $k$ and $m$, as described in the DPR process in Section 5.1.
+These trained GNN models are denoted as GNN $(k,m)$.
+Additional implementation details related to this section are available in Appendix C.
+As illustrated in the example from our test set (Figure 3), the DPR-based approach fails to retrieve a key passage related to the main contributions of the paper, whereas our GNN successfully identifies it.
+Therefore, graph-based representation, which leverages both sequential and hierarchical node connections, effectively identifies the salient passages required to enhance the feedback generation process.
+The GNN extracted passages, along with a structured prompt, serve as the input in the LLM fine-tuning stage, as discussed in Section 5.5.
+Figure 3: Excerpt of a test sample where passages highlighted in cyan colour refer to relevant passages identified by GNN only and yellow colour refer to relevant passages identified by both GNN and DPR
+5.4 Grounding and Inference Efficiency
+For practical deployment, our GNN-based method offers a crucial advantage by operating autonomously at inference time.
+Unlike DPR-based approaches, which rely on human-generated feedback to act as queries, that may be unavailable in real-world scenarios, our system identifies salient passages directly from the document graph, thereby eliminating the dependency on external grounding.
+Unlike SEA baselines that process the full text, AutoRev generates feedback from a reduced set of GNN-selected passages.
+This approach significantly lowers computational cost and input length.
+As noted by Hsieh et al.
+(2024), shorter contexts enhance LLM performance, which is also reflected in our higher scores compared to baselines (Section 6).
+Table 2 details the specific reductions in tokens and passages.
+Method Avg Token Len Avg Passages
+Full Paper 9445.92 47.07
+GNN $(5,5)$ 4147.15 29.49
+GNN $(5,3)$ 2899.59 21.00
+Table 2: Average token length (Avg Token Len) and Average passage count (Avg. Passages) for a full paper, our best-configuration: GNN $(5,5)$, and our fastest-configuration: GNN $(5,3)$.
+5.5 Fine-Tuning Large Language Models for Structured Feedback Generation
+General-purpose LLMs are not trained specifically to generate structured feedback for academic research papers.
+Our motivation for fine-tuning is to ensure that the models learn to generate qualitatively sound, factual and structured feedback.
+We select Mistral-Instruct-v0.2 (Mistral) Jiang et al.
+(2023) to align with our SEA-EA baseline, and state-of-the-art (SOTA), open-weight model, Llama 3.1 8B (Llama) Grattafiori et al.
+(2024) for its strong text generation capabilities.
+To efficiently adapt these models for structured feedback generation while ensuring computational feasibility, we employ Low-Rank Adaptation (LoRA) Hu et al.
+(2022) for finetuning.
+We use the trained GNN, as described in Section 5.3, to retrieve relevant passages from each paper, which are provided as input to the LLM during fine-tuning.
+For supervision, we use the Summary, Strengths, Weaknesses, and Questions sections of human-written peer reviews for ICLR 2024 papers, consolidated using GPT-4 and released as part of the dataset introduced by Yu et al.
+(2024).
+These consolidated reviews serve as structured feedback targets for training the LLM.
+The dataset split is summarized in Table 3.
+Dataset Split Number of Samples
+Training Set 1,765
+Test Set 271
+Table 3: Number of samples used for training and testing sourced from Yu et al. (2024)
+Both models are fine-tuned on passages retrieved via varying GNN $(k,m)$ configurations (Section 5.3), denoted as Mistral $(k,m)$ and Llama $(k,m)$.
+Prompts and hyperparameters are detailed in Appendix D.
+6 Results and Analysis
+Model ROUGE F1 ROUGE Recall BERTScore-F1 Training Hours
+R-1 R-2 R-L R-1 R-2 R-L
+M-7B 22.55 8.89 11.28 13.48 5.29 16.73 83.79
+SEA-E 34.29 11.71 14.98 23.28 7.90 10.13 84.04
+SEA-EA 36.97 13.02 15.88 25.62 8.97 10.97 84.15
+Mistral $(5,3)$ 53.44 15.47 22.39 53.18 15.42 22.31 86.44 4.34
+Llama $(5,3)$ 54.24 15.91 22.59 53.75 15.77 22.41 86.59 3.60
+Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
+Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
+Table 4: Performance on the ICLR-24 test set comparing our models with baselines (M-7B, SEA-E, SEA-EA) from Yu et al.
+(2024).
+Bold and underlined values indicate best and second-best scores, respectively.
+Please note that the GPU configurations differ between Llama and Mistral fine-tuning (details in Appendix D).
+This section presents a comprehensive evaluation of AutoRev across automatic, human, and robustness-focused analyses.
+We first benchmark feedback quality on the ICLR 2024 test set using ROUGE-1/2/L (Recall and F1) Lin (2004) and BERTScore-F1 Zhang* et al.
+(2020) against the sections in the reference feedback, assessing both lexical overlap and semantic similarity.
+To evaluate feedback quality beyond surface similarity, we additionally report Aspect Coverage (ACov) Yuan et al.
+(2022), which measures the breadth of scholarly aspects addressed in the generated feedback.
+Given the limitations of automated metrics, we conduct human evaluations assessing qualitative dimensions such as feedback depth, constructiveness, thoroughness, and helpfulness, alongside a distinct assessment of factual validity of the generated feedback.
+We further examine cross-venue generalization on NeurIPS and COLM 2025 papers and assess practical utility via an impact-oriented user survey.
+Finally, we employ a scalable LLM-as-a-Judge framework and perform ablation studies to analyze the effects of passage-selection parameters.
+6.1 Performance Comparison with Baselines
+We compare AutoRev against SEA-based baselines from Yu et al.
+(2024), including M-7B, SEA-E, and SEA-EA.
+All baselines operate on the full paper text during both training and inference, without any learned passage selection.
+M-7B performs direct inference using an instruction-tuned Mistral model, SEA-E leverages standardized fine-tuning to improve review quality, and SEA-EA further incorporates a mismatch score to assess consistency between paper content and generated review.
+In contrast, AutoRev employs learned passage selection using a Graph Neural Network over a document graph, providing only salient passages to the LLM.
+As shown in Table 4 , AutoRev consistently outperforms all baselines across ROUGE (Recall and F1) and BERTScore metrics.
+Our best-performing configuration, Llama $(5,5)$ , achieves an average improvement of 58.72% over the state-of-the-art baselines.
+To isolate the effects of model capacity from retrieval quality, we also evaluate AutoRev using Mistral.
+The substantial gains observed with Mistral confirm that the performance improvements stem primarily from GNN-based passage selection rather than the choice of LLM.
+The graph-based retrieval enables the model to capture both hierarchical and sequential dependencies across the document, providing the focused context necessary for generating comprehensive feedback.
+Finally, Llama $(5,3)$ offers a strong efficiency-performance trade-off, reducing training time by 32.07% with only a 0.71% decrease in performance relative to Llama $(5,5)$ , while still outperforming all baselines.
+6.2 Aspect Coverage Analysis
+We evaluate the breadth of scholarly criteria addressed in the generated structured feedback using Aspect Coverage (ACov), following the aspect tagging framework of Yuan et al.
+(2022).
+Unlike surface-level similarity metrics, ACov measures whether feedback engages with the diverse dimensions typically considered in peer review.
+Table 5 reports aspect coverage for our two best-performing configurations, Llama $(5,3)$ and Llama $(5,5)$.
+Both variants demonstrate consistently strong coverage across core feedback dimensions, except for Motivation/Impact and Replicability, indicating that AutoRev generates feedback that goes beyond summarization and engages with substantive and technical aspects of the paper.
+Overall, these results suggest that AutoRev provides broad and balanced coverage of evaluation criteria, reinforcing its suitability as a comprehensive pre-submission feedback system.
+Model SUM CLA SOU SUB ORI CMP REP MOT
+Llama $(5,3)$ 100.0 99.6 98.9 95.2 92.6 79.0 72.0 66.8
+Llama $(5,5)$ 100.0 99.3 99.3 95.9 90.0 84.5 71.6 66.8
+Table 5: Aspect Coverage (%) of model-generated structured feedback evaluated using the Aspect Tagger (Yuan et al., 2022) across aspects: Summary (SUM), Clarity (CLA), Soundness/Correctness (SOU), Substance (SUB), Originality (ORI), Meaningful Comparison (CMP), Replicability (REP), and Motivation/Impact (MOT).
+6.3 Human Evaluation
+Understanding the limitations of automatic evaluation metrics, as highlighted by Hanna and Bojar (2021) and Schluter (2017), we conduct human evaluations to assess the acceptability, factual grounding, and practical utility of the generated feedback.
+Our evaluation is structured around two complementary dimensions: (i) qualitative assessment of feedback quality and (ii) factual validity of individual claims.
+We perform these evaluations on both the ICLR 2024 test set and a small out-of-domain dataset drawn from NeurIPS and COLM 2025.
+Due to the cost of expert annotation, human evaluations are conducted on representative subsets of the data.
+We report results for Llama $(5,3)$ and Llama $(5,5)$, our best-performing configurations.
+6.3.1 Qualitative Assessment
+A total of 30 papers are randomly selected from the ICLR 2024 test set (ICT_30).
+Each paper is associated with three feedback instances: one reference feedback corresponding to the ICT_30 paper and one generated by each of Llama $(5,3)$ and Llama $(5,5)$, yielding 90 feedback samples.
+A total of 27 annotators participate in the study, with each feedback annotated by three independent annotators.
+Feedback quality is evaluated across four criteria:
+  1. 1.
+    Feedback Depth (FD): Evaluates the technical understanding and meaningful engagement with the paper’s content.
+  2. 2.
+    Feedback Thoroughness (FT): Measures the coverage of key components like methodology, results, and reproducibility.
+  3. 3.
+    Feedback Constructiveness (FC): Assesses the utility of suggestions for improvement.
+  4. 4.
+    Feedback Helpfulness (FH): Evaluates how effectively important issues are identified and explained to the authors.
+Annotators assign scores on a 1-5 scale following the guidelines detailed in Appendix E and snapshots of the annotation interface are shown in Appendix F.
+Scores are averaged across annotators for each feedback and then aggregated across all samples for each model.
+As shown in Table 6, AutoRev-generated feedback closely approaches ICT_30 feedback in constructiveness and helpfulness, which are particularly critical for pre-submission support.
+While ICT_30 feedback remains superior in depth and thoroughness, the results show that AutoRev does not merely summarize paper content, but meaningfully engages with methodological and clarity issues, reinforcing its suitability as a structured feedback tool for authors.
+Feedback FD FT FC FH
+ICT_30 3.54 3.49 3.26 3.41
+Llama $(5,3)$ 3.37 3.30 3.27 3.34
+Llama $(5,5)$ 3.27 3.24 3.19 3.30
+Table 6: Human Qualitative Assessment scores (scale of 1–5) of generated feedback using Llama models against reference feedback corresponding to the ICT_30 paper across four criteria: Feedback Depth (FD), Feedback Thoroughness (FT), Feedback Constructiveness (FC), and Feedback Helpfulness (FH).
+6.3.2 Factual Validity Assessment
+To evaluate factual grounding, we randomly select 10 papers from the ICLR 2024 test set (ICT_10).
+Each paper is associated with three feedback instances: one reference feedback corresponding to the ICT_10 paper and outputs from Llama $(5,3)$ and Llama $(5,5)$.
+These feedbacks are annotated by 10 annotators, with each paper evaluated by two annotators.
+Annotators assess each statement within a feedback based on its factual consistency with the source paper, using a four-point scale: Completely Agree, Mostly Agree, Mostly Disagree, and Completely Disagree.
+The evaluation guidelines are provided in Appendix G, and snapshots of the annotation interface are shown in Appendix F.
+The percentage agreement scores summarized in Tables 7 and 12 indicate that AutoRev maintains strong factual grounding in the source document.
+While ICT_10 feedback achieves the highest agreement, AutoRev-generated feedback receives a clear majority of positive ratings, with complete disagreements (potential hallucinations) remaining below 15%.
+These findings suggest that graph-based passage selection effectively anchors feedback generation in relevant evidence, mitigating hallucinated content.
+6.3.3 Cross-Venue Generalization Assessment
+To further examine robustness beyond the training distribution, we conduct an out-of-domain human evaluation on a curated set of 10 papers, comprising five papers each from COLM 2025 and NeurIPS 2025 (CNT_10).
+Each paper is evaluated using feedback generated by Llama $(5,3)$ and Llama $(5,5)$.
+Due to the absence of publicly available consolidated reference feedback for these venues and the high cost of expert annotation, reference feedback is not included.
+A total of 10 annotators participate in the study, with each paper evaluated by two annotators following the same guidelines described in Section 6.3.2.
+The results, reported in Table 7 and 12, show comparable factual agreement across venues.
+This suggests that AutoRev generalizes beyond the ICLR 2024 training distribution, capturing structural and semantic regularities common to scientific writing rather than overfitting to venue-specific conventions.
+Dataset Feedback CA MA MD CD
+ICT_10 Reference 39.12 37.85 16.40 6.62
+Llama $(5,3)$ 32.64 31.94 22.92 12.50
+Llama $(5,5)$ 32.27 32.98 23.05 11.70
+CNT_10 Llama $(5,3)$ 22.00 39.86 24.57 13.57
+Llama $(5,5)$ 24.06 40.25 22.19 13.50
+Table 7: Factual validity assessment (% agreement) on ICLR 2024 (ICT_10) and cross-venue CNT_10 datasets. CA/MA indicate grounded statements, while CD reflects clear factual errors (hallucinations).
+6.3.4 Impact Assessment
+To evaluate the impact of our system, we conduct a survey among the annotators involved in the tasks discussed in Sections 6.3.2 and 6.3.3.
+Participants are asked: “As an author, how likely are you to use AutoRev to get feedback for your own manuscript before submitting it to a conference?” On a 5-point scale, the system received an average rating of 4.2.
+This indicates that authors are strongly inclined to use our model for pre-submission critique, demonstrating the system’s practical utility.
+6.4 LLM-as-a-Judge Evaluation
+As mentioned by Zheng et al.
+(2023), the LLM-as-a-Judge method offers an explainable and affordable way to approximate human preferences, which are otherwise costly to gather.
+We employ a large language model, Gemini-2.0-Flash [^3] https://deepmind.google/technologies/gemini/ as an automated judge to evaluate generated feedback.
+Since LLM-as-a-Judge evaluation is scalable and much cheaper to undertake, we evaluate all our models against the reference feedback, over the entire ICT dataset.
+The evaluation criteria and prompting guidelines are adapted from those used in our human qualitative assessment (Section 6.3.1; Appendix E).
+The results of this experiment are summarized in Table 8
+A key finding is the strong correlation between LLM-as-a-Judge scores and human annotations (Table 6).
+This alignment validates the automated judge as a reliable proxy for assessing feedback quality in this domain.
+Feedback FD FT FC FH
+ICT 3.98 3.87 3.79 3.79
+Mistral $(5,3)$ 3.50 3.35 3.44 3.32
+Llama $(5,3)$ 3.61 3.45 3.47 3.39
+Mistral $(5,5)$ 3.48 3.36 3.43 3.34
+Llama $(5,5)$ 3.63 3.44 3.47 3.41
+Table 8: LLM-as-a-Judge (Gemini-2.0-Flash) scores (scale of 1–5) comparing model-generated feedback against ICT across Feedback Depth (FD), Feedback Thoroughness (FT), Feedback Constructiveness (FC), and Feedback Helpfulness (FH).
+Bold and underlined values indicate the best and second-best model scores (excluding reference feedback), respectively.
+6.5 Ablation Studies
+We perform ablation studies by systematically varying the values of $k\in\{1,3,5\}$ and $m\in\{3,5\}$ during LLM fine-tuning.
+The detailed results are reported in Appendix B.
+The results indicate that increasing $k$ and $m$ provides richer context to both Mistral and Llama to generate feedback, leading to consistent improvements across all evaluation metrics.
+Though, this comes at the downside of higher training times as $m$ increases for a fixed $k$.
+7 Conclusion
+In this work, we present AutoRev, an author-centric system for generating structured pre-submission feedback using graph-based passage retrieval.
+By modeling papers as hierarchical document graphs and leveraging a GNN to identify salient content, AutoRev substantially reduces LLM input length, overcoming the limitations of full-paper fine-tuning, stemming from long-context degradation Hsieh et al.
+(2024), while operating without human-written grounding.
+Experiments on ICLR 2024 show that AutoRev demonstrates strong empirical performance through automatic metrics, with human evaluations confirming constructive, well-grounded feedback and cross-venue generalization.
+Notably, Llama $(5,5)$ excels in generating comprehensive feedback for rigorous refinement of manuscripts, whereas Llama $(5,3)$ offers a favorable efficiency-performance trade-off for large-scale deployment.
+Limitations
+While our framework provides a robust and high-performing solution for feedback generation, we acknowledge that no automatic system can fully replicate human judgment in complex tasks such as peer reviewing, as noted by Ye et al.
+(2024).
+Our goal is to assist authors in refining their papers and we aim for our system to support authors only in the pre-submission feedback phase.
+Currently, our system is primarily focused on ICLR 2024 papers due to the limited availability of public datasets containing consolidated reviews.
+However, the consistent positive performance observed in out-of-domain human evaluations (Section 6.3.3) suggests that our model generalizes well beyond the ICLR 2024 setting.
+Another limitation is our reliance on human-written reviews consolidated using GPT-4.
+While consolidated reviews provide a clearer and more consistent supervision signal than multi-review datasets, which are often noisy and contradictory and can hinder model learning, this approach introduces its own challenges.
+In particular, relying on an LLM for consolidation may induce biases.
+Future work can explore hybrid strategies that reduce or eliminate reliance on LLMs for consolidating multiple reviews into a single supervision signal.
+Ethical Considerations
+The deployment of AutoRev raises ethical concerns regarding bias, misuse, and academic integrity.
+Critically, while the system is designed strictly as an author-centric aid, there is a risk of over-reliance or misuse in decision-making, which could undermine the nuances of human reviewing.
+Furthermore, by training on historical peer reviews and GPT-4 consolidated data, the model may inherit and amplify existing biases present in those texts, such as leniency toward well-known authors or under-representation of certain research areas, potentially skewing evaluations.
+Additionally, the high fidelity of the generated text raises concerns about academic misconduct, where AI critiques could be misrepresented as original human work.
+To address these issues, we advocate for transparency and emphasize that AutoRev is intended solely as a supplementary tool to support, not replace, the scientific review process.
+Appendix A Graph Construction
+Formally, each paper is represented as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where nodes $\mathcal{V}$ represent text segments with semantically encoded features, and $\mathcal{E}$ denote the hierarchical and sequential relationships that form the document’s structure.
 Our graph construction is designed to be generalizable across diverse academic paper structures, accommodating arbitrary levels of nesting and heterogeneous section organization.
-4.2.1. Node Definitions
+A.1 Node Definitions
   • Paper Node ($v_{\text{paper}}$): A single root node representing the entire document.
@@
 […]
-  • Sentence Nodes ($\mathcal{T}_{jk}=\{t_{jk1},t_{jk2},\dots,t_{jkp}\}$): Nodes for individual sentences within each passage $o_{jk}$, tokenized using NLTK (Bird and Loper, 2004).
-4.2.2. Edge Definitions
+  • Sentence Nodes ($\mathcal{T}_{jk}=\{t_{jk1},t_{jk2},\dots,t_{jkp}\}$): Nodes for individual sentences within each passage $o_{jk}$, tokenized using NLTK Bird and Loper (2004).
+A.2 Edge Definitions
 Edges capture both hierarchical and sequential relationships within the document.
@@
 The final edge set is the union of hierarchical and sequential edges:
-(1) $\mathcal{E}=\mathcal{E}_{\text{hier}}\cup\mathcal{E}_{\text{seq}},$
+$\mathcal{E}=\mathcal{E}_{\text{hier}}\cup\mathcal{E}_{\text{seq}},$ (1)
 where $\mathcal{E}_{\text{hier}}$ captures the vertical (structural) relationships and $\mathcal{E}_{\text{seq}}$ captures the horizontal (textual progression) flow.
@@
 A toy example of such a graph, built from a paper with 2 Headings, 2 Sub-Headings, 4 Passages, and 6 Sentences, is shown in Figure 1.
-4.3. GNN Training for Multi Modal Passage Retrieval
-Figure 3.
-Excerpt of a test sample where passages highlighted in cyan colour refer to relevant passages identified by GNN only and yellow colour refer to relevant passages identified by both GNN and DPR:
-To train a GNN on the graph constructed in Section 4.2, we take the data generated in Section 4.1, which maps each paper $p_{i}$ to a set of important passages, $\mathcal{O}_{i}=\{o_{1},o_{2},\dots,o_{n}\}$.
-Each passage is important for generating a review for the paper it is associated with.
-This mapping is used as a training set for training the GNN.
-The node embeddings for sentence nodes are initialized with embeddings from a Sentence-Transformer (Reimers and Gurevych, 2019) model, namely all-mpnet-base-v2.
-The node embeddings for all other nodes are initialized with zero embeddings of dimension $768$, matching that of the Sentence-Transformer model.
-Now, a Graph Attention Network (Veličković et al., 2018) is used to train the graph to predict the important passages.
-After passing the initial representation of the graph nodes through three layers of GATConv (PyTorch Geometric’s implementation of Graph Attention Networks) [^6] https://pytorch-geometric.readthedocs.io/en/latest/generated/torch_geometric.nn.conv.GATConv.html, we apply a non-linear classifier to the final passage node representations.
-The classification is defined as follows:
-  1. (1)
-    A label of 0 indicates that the passage is not important for review generation.
-  2. (2)
-    A label of 1 indicates that the passage is important for review generation.
-The GNN is trained using various configurations based on the values of $k$ and $m$, as described in the DPR process in Section 4.1. These trained GNN models are denoted as GNN $(k,m)$.
-We hypothesize that our choice of graph construction which models the semantic representation of nodes and flow of information through structure, allows our model to capture the intricate inter-dependencies of sentences and passages across different sections within an academic paper, and effectively make a choice about which passage(s) is/are important.
-The dependencies, according to our hypothesis, are learned from the sequential and hierarchical connections that we introduce in our graph construction method.
-The passing of information from one node to another is enhanced by the GATConv (Equations 2, 3, and 4) layers, which involves the concept of attention to selectively propagate messages in the Message Passing layer of a GNN.
-(2) $\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$
-(3) $e_{ij}=\text{LeakyReLU}\Bigl(\mathbf{a}^{T}\bigl[\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}\bigr]\Bigr)$
-(4) $\alpha_{ij}=\frac{\exp\bigl(e_{ij}\bigr)}{\sum_{k\in\mathcal{N}(i)}\exp\bigl(e_{ik}\bigr)}$
-Additional implementation details related to this section are available in Appendix B.
-As illustrated in the example from our test set (Figure 3), the DPR-based approach fails to retrieve a key passage related to the main contributions of the paper, whereas our GNN successfully identifies it.
-Therefore, our multi-modal graph, which leverages both sequential and hierarchical node connections, more effectively identifies the salient passages required to enhance the review generation process.
-The GNN extracted passages, along with a structured prompt, serve as the input in the LLM fine-tuning stage, as discussed in Section 4.5.
-4.4. Grounding Requirements and Inference Efficiency
-For practical deployment, our GNN-based method offers a crucial advantage by operating autonomously at inference time.
-It identifies important passages directly from the paper’s graph structure, eliminating the need for grounding from human-generated reviews which act as queries in the DPR-based approaches, but may be unavailable in real-world scenarios.
-Second, by selecting only the passages deemed most relevant, we significantly reduce the input length and associated computational costs for downstream LLM-based review generation.
-This not only leads to higher performance of LLMs, as outlined by Hsieh et al.
-(2024) but also leads to higher scores of the generated reviews compared to our baselines (see Section 5).
-The statistics related to the decrease in number of passages and token length post-extraction by the GNN model are reported in Table 2.
-Table 2. Average token length (Avg. Token Len) and Average passage count (Avg. Passages) for a full paper, our best-configuration: GNN $(5,5)$, and our fastest-configuration: GNN $(5,3)$.
-Input to LLM Avg Token Len Avg Passages
-Full Paper 9445.92 47.07
-GNN $(5,5)$ 4147.15 29.49
-GNN $(5,3)$ 2899.59 21.00
-4.5. Fine-Tuning Large Language Models for Review Generation
-General-purpose LLMs are not trained specifically to generate peer reviews for academic research papers.
-Our motivation for fine-tuning is to ensure that the models learn to generate qualitatively sound, factual and structured peer reviews.
-We select Mistral-Instruct-v0.2 (Mistral) [^7] https://arxiv.org/abs/2310.06825, inspired by our baseline SEA-EA, which also uses Mistral-Instruct-v0.2 as its underlying backbone model.
-We also use Llama 3.1 8B (Llama) [^8] https://arxiv.org/abs/2407.21783, as it is an open-weight state-of-the-art (SOTA) model with strong text generation capabilities for a variety of tasks.
-We employ LoRA (Low-Rank Adaptation) (Hu et al., 2022), for fine-tuning due to its efficiency in adapting large language models to particular tasks while maintaining computational feasibility.
-This allows us to leverage the pre-trained capabilities of Llama and Mistral while efficiently adapting them for review generation.
-We use the trained GNN, as described in Section 4.3, to retrieve relevant passages for each paper that serve as the LLM’s input during fine-tuning.
-The human-written peer reviews for ICLR 2024 papers, which are consolidated with GPT-4, served as the ground truth for fine-tuning the LLM used to generate the final reviews.
-These reviews are released as a part of the dataset released by Yu et al.
-(2024).
-The dataset split is summarized in Table 3.
-Table 3. Number of samples used for training and testing sourced from Yu et al. (2024)
-Dataset Split Number of Samples
-Training Set 1,765
-Test Set 271
-Both Mistral and Llama are fine-tuned on the passages retrieved, using various configurations of the trained GNN $(k,m)$, as described in Section 4.3.
-These fine-tuned models are denoted Mistral $(k,m)$ and Llama $(k,m)$, respectively.
-The prompt and hyperparameter settings used during fine-tuning are provided in Appendix C.
-5. Results and Analysis
-Table 4.
-Overall performance on the ICLR-24 reviews test dataset, comparing M-7B, SEA-E, SEA-EA as baselines from Yu et al.
-(2024) with our fine-tuned models.
-For both Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral), the configuration with $(5,5)$ achieves the best performance, while $(5,3)$ has the shortest training time.
-Bold highlights the best scores, and underlined values indicate the second-best.
-Please note that the GPU configurations for finetuning Llama and Mistral were different.
-For details, refer to Appendix C.
-Model ROUGE F1 ROUGE Recall BERTScore-F1 Training Time (hrs)
-R-1 R-2 R-L R-1 R-2 R-L
-M-7B 22.55 8.89 11.28 13.48 5.29 16.73 83.79
-SEA-E 34.29 11.71 14.98 23.28 7.90 10.13 84.04
-SEA-EA 36.97 13.02 15.88 25.62 8.97 10.97 84.15
-Mistral $(5,3)$ 53.44 15.47 22.39 53.18 15.42 22.31 86.44 4.34
-Llama $(5,3)$ 54.24 15.91 22.59 53.75 15.77 22.41 86.59 3.60
-Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
-Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
-This section details our system’s performance analysis.
-Section 5.1 reports the evaluation of generated review quality using ROUGE-1, ROUGE-2, and ROUGE-L (Recall and F1) (Lin, 2004), as well as BERTScore-F1 (Zhang* et al., 2020), on the GPT-4-consolidated reviews test set for ICLR 2024 papers.
-ROUGE measures n-gram overlap, with recall reflecting content coverage and F1 balancing precision and recall, while BERTScore captures semantic similarity in embedding space.
-Section 5.2 further employs the Aspect Coverage (ACov) metric (Yuan et al., 2022), designed specifically for assessing peer-review quality.
-Given the limitations of automated metrics, we also conduct human evaluations (Section 5.3) to assess review acceptability.
-To address the inherent scalability challenges of manual assessment, we employ an LLM-as-a-Judge framework (Section 5.4), which shows strong correlation with human judgments.
-5.1. Performance Comparison with Baselines
-We compare the performance of our models against baselines sourced from Yu et al.
-(2024).
-As reported in Table 4, the M-7B baseline performs direct inference with Mistral guided by an instruction prompt to generate peer reviews, SEA-E baseline leverages standardized fine-tuning data to produce more constructive reviews, while the SEA-EA baseline further introduces a mismatch score to evaluate the consistency between paper content and generated reviews.
-Our results suggest that we consistently outperform the baselines on all metrics, with significant improvements in the ROUGE metrics (Recall and F1).
-Our best-performing configuration, Llama $(5,5)$ (selected based on the highest BERTScore), achieves an improvement of 58.72% over the state-of-the-art (SOTA) baselines across all evaluation metrics.
-To ensure that these gains are not solely due to using a more capable LLM, we also evaluate our method using Mistral.
-The results demonstrate that even with Mistral, integrating our Multi Modal GNN-based passage selection method leads to substantial improvements, confirming that the performance boost is primarily due to our graph-based retrieval strategy rather than the choice of the LLM alone.
-The integration of a Graph Attention Network (GAT) enables the model to effectively capture complex dependencies between relevant passages and sentences within the paper, offering the essential context required to generate comprehensive reviews.
-Our fastest, second-best, fine-tuned model, Llama $(5,3)$, performs competitively while being highly efficient.
-It reduces the training time by 32.07% with only a 0.71% drop in performance compared to our best model, consistently outperforming all baselines.
-This shows our GNN-based approach achieves near-optimal performance with significantly less computational overhead.
-5.2. Aspect Coverage Analysis
-As discussed by Yuan et al.
-(2022), each comment in a peer review typically addresses diverse aspects of an academic research paper, such as Summary (SUM), Motivation/Impact (MOT), Originality (ORI), Soundness/Correctness (SOU), Substance (SUB), Replicability (REP), Meaningful Comparison (CMP), and Clarity (CLA).
-These cannot be reliably assessed by conventional automatic metrics like ROUGE-F1 or BERTScore.
-To address this, Yuan et al.
-(2022) introduces an Aspect Tagger which can be used to compute Aspect Coverage (ACov) that measures the proportion of aspects covered by a peer-review.
-We evaluate the two best-performing models Llama $(5,5)$ and Llama $(5,3)$ using this tagger.
-Table 5.
-Aspect Coverage (%) of model-generated reviews evaluated using the Aspect Tagger (Yuan et al., 2022).
-Summary (SUM), Clarity (CLA), Soundness/Correctness (SOU), Substance (SUB), Originality (ORI), Meaningful Comparison (CMP), Replicability (REP), and Motivation/Impact (MOT).
-Model SUM CLA SOU SUB ORI CMP REP MOT
-Llama (5,3) 100.0 99.6 98.9 95.2 92.6 79.0 72.0 66.8
-Llama (5,5) 100.0 99.3 99.3 95.9 90.0 84.5 71.6 66.8
-As shown in Table 5, both variants achieve strong coverage for Summary, Clarity, and Soundness, indicating the reliability of our model-generated reviews in capturing these core aspects.
-However, the aspect coverage for Motivation/Impact and Replicability is comparatively lower.
-Overall, the results suggest that our model captures most key aspects, indicating that the generated reviews are comprehensive and address the quality of multiple dimensions of a research paper.
-5.3. Human Evaluation
-Understanding the limitations of automatic evaluation metrics as highlighted by Hanna and Bojar (2021) and Schluter (2017), we conduct human evaluations to assess the acceptability of the generated reviews.
-We model our evaluation around two distinct aspects of review quality.
-The first is a Qualitative Assessment, where the generated review is evaluated in isolation across four key criteria: Reviewer Confidence, Review Thoroughness, Review Constructiveness, and Review Helpfulness.
-The second is a Factual Validity Assessment, which ensures the review’s specific claims are factually grounded in the source document.
-Since human evaluation is very costly, we conduct each of the two tasks on a subset of the test dataset, as described in their individual subsections.
-Additionally, we only considered Llama $(5,3)$ and Llama $(5,5)$, our best performing models for the evaluation process.
-5.3.1. Qualitative Assessment
-A total of 30 papers are chosen at random from our test set.
-Each of these papers are associated with 3 reviews, one Gold, one from Llama $(5,3)$, and one from Llama $(5,5)$, thus, resulting in a total set of 90 reviews.
-A total of 27 annotators are chosen, and each annotator is given 10 reviews to annotate.
-The evaluation criteria is designed to capture key aspects of review quality and include:
-  1. (1)
-    Reviewer Confidence (RCf): Reflects the reviews’ depth of technical understanding and engagement with the content.
-  2. (2)
-    Review Thoroughness (RT): Measures how well the review covers the methodology, results, and reproducibility.
-  3. (3)
-    Review Constructiveness (RCt): Assesses the clarity and usefulness of suggestions for improvement.
-  4. (4)
-    Review Helpfulness (RH): Indicates how well the review identifies important issues and explains them clearly to the authors.
-Each annotator is tasked to assign a score from 1 to 5 for each of the above criteria.
-The evaluation guidelines are detailed in Appendix D.
-We end up with each review being annotated by 3 annotators.
-Score for a review is averaged across all its annotators, and the final score of a model, is the average score across all its reviews.
-The results of this experiment are summarized in Table 6.
-Table 6.
-Results of our human qualitative assessment.
-This table shows average scores (on a scale of 1-5) comparing gold-standard reviews with those generated by our two best-performing Llama models across four key criteria: Reviewer Confidence (RCf), Review Thoroughness (RT), Review Constructiveness (RCt), and Review Helpfulness (RH).
-Review RCf RT RCt RH
-Gold 3.54 3.49 3.26 3.41
-Llama (5,3) 3.37 3.30 3.27 3.34
-Llama (5,5) 3.27 3.24 3.19 3.30
-5.3.2. Factual Validity Assessment
-A total of 10 papers are chosen at random from our test set.
-Each of these papers are associated with 3 reviews, one Gold, and two generated by our best performing models.
-Thus, the final evaluation set consists of 10 papers, each associated with 3 reviews, one Gold, one from Llama $(5,3)$, and one from Llama $(5,5)$.
-These 10 papers are circulated among 10 annotators.
-It is ensured that a single paper is annotated by at least 2 annotators.
-Annotators are instructed to evaluate each statement within a review based on its factual grounding in the full source paper.
-For each point, they assign a rating from a four-point scale: ‘Completely Agree’, ‘Mostly Agree’, ‘Mostly Disagree’, or ‘Completely Disagree’.
-The evaluation guidelines are detailed in Appendix D.
-The results of this experiment are summarized in Table 7 and Table 8.
-Table 7.
-Detailed results of the Factual Validity Assessment.
-This table presents the percentage distribution of human agreement ratings for claims made in reviews from our Llama models and the gold standard.
-The abbreviations are: Completely Agree (CA), Mostly Agree (MA), Mostly Disagree (MD), and Completely Disagree (CD).
-Review CA MA MD CD
-Gold 39.12% 37.85% 16.40% 6.62%
-Llama (5,3) 32.64% 31.94% 22.92% 12.50%
-Llama (5,5) 32.27% 32.98% 23.05% 11.70%
-Table 8.
-Detailed results of the Factual Validity Assessment.
-This table shows the overall Positive Rating Percentage, calculated by summing the “Completely Agree” and “Mostly Agree” categories from Table 7, which indicates the factual soundness of the reviews.
-Review Positive Rating %
-Gold 76.97%
-Llama (5,3) 65.25%
-Llama (5,5) 64.58%
-5.4. LLM-as-a-Judge Evaluation
-As mentioned by Zheng et al.
-(2023), the LLM-as-a-Judge method offers an explainable and affordable way to approximate human preferences, which are otherwise costly to gather.
-We employ a large language model, Gemini-2.0-Flash [^9] https://deepmind.google/technologies/gemini/ as an automated judge to evaluate the quality of reviews generated by our proposed model as compared to the gold-standard reviews.
-For the LLM-as-a-judge task, the evaluation criteria and prompting guidelines are adapted from those used in our human qualitative assessment, which are detailed in Section 5.3.1.
-Since LLM-as-a-Judge evaluation is scalable and much cheaper to undertake, we evaluate all our models against the gold review, over the entire test dataset.
-A key finding from our evaluation is the strong correlation between the LLM-as-a-Judge scores (Table 9) and the results from our human annotators (Table 6).
-This alignment suggests that the LLM-as-a-Judge is a reliable and valid proxy for assessing review quality in this domain.
-Table 9.
-Qualitative evaluation of model-generated reviews using Gemini-2.0-Flash as the LLM judge.
-The table presents the average scores (out of 5) for all model configurations and the gold standard across four criteria: Reviewer Confidence (RCf), Review Thoroughness (RT), Review Constructiveness (RCt), and Review Helpfulness (RH).
-Bold values indicate the best score in each column (excluding Gold), and underlined values indicate the second-best.
-Review RCf RT RCt RH
-Gold 3.98 3.87 3.79 3.79
-Mistral $(5,3)$ 3.50 3.35 3.44 3.32
-Llama $(5,3)$ 3.61 3.45 3.47 3.39
-Mistral $(5,5)$ 3.48 3.36 3.43 3.34
-Llama $(5,5)$ 3.63 3.44 3.47 3.41
-5.5. Ablation Studies
-We perform ablation studies by systematically varying the values of $k\in\{1,3,5\}$ and $m\in\{3,5\}$ during LLM fine-tuning.
-The detailed results are reported in Appendix A.
-The results indicate that increasing $k$ and $m$ provides richer context to both Mistral and Llama to generate reviews, leading to consistent improvements across all evaluation metrics.
-Additionally, we observe that higher values of $m$ result in longer training times, as the GNN retrieves and passes a larger set of passages as input for finetuning the LLMs.
-6. Limitations
-While our framework provides a robust and high-performing solution for review generation, we acknowledge that no automatic system can fully replicate human judgment in complex tasks such as peer reviewing, as noted by Ye et al.
-(2024).
-Our goal is to assist authors in refining their papers and we aim for our system to support human reviewers.
-Currently, our system is primarily focused on ICLR 2024 papers due to the limited availability of public datasets containing consolidated reviews.
-However, we believe in the generalization capabilities of our model, given the proven effectiveness of graph-based approaches in various downstream NLP tasks, as discussed in Section 1.
-Our reliance on human written reviews, consolidated using GPT-4 represents another limitation.
-While multi-review datasets more closely reflect the diversity of real peer review, they are often noisy and contradictory, which can hinder model learning.
-Consolidated reviews provide a clearer and more consistent supervision signal, but at the cost of reducing the richness of differing reviewer perspectives.
-Future work should explore hybrid approaches that retain both consistency and diversity.
-7. Ethical Considerations
-The deployment of AutoRev raises key ethical concerns related to bias, misuse, and academic integrity.
-Since the system is trained on existing peer reviews, it may inherit and amplify biases present in those texts, such as leniency toward well-known authors or under-representation of certain research areas, potentially skewing evaluations.
-Moreover, there is a risk that users may over-rely on AutoRev for making or substituting final review decisions, which could undermine the depth and critical thinking central to human peer review.
-Additionally, the realistic quality of AI-generated reviews opens the possibility of unethical use, such as submitting AI-authored critiques as original human work.
-To address these issues, we advocate for responsible usage of AutoRev as a supplementary tool, accompanied by transparency in AI involvement and safeguards to preserve academic integrity.
-8. Conclusion
-In this work, we introduce AutoRev, an automatic peer review generation system designed specifically for academic research papers.
-Our approach leverages a novel graph-based framework to represent the structure of a paper which enables the effective extraction of important passages using a Graph Neural Network.
-This strategy of selecting passages and reducing the input length to LLMs addresses the limitations of traditional methods, like finetuning using the entire paper as input (Yu et al., 2024), which leads to sub-optimal performance of LLMs due to large input token length (Hsieh et al., 2024).
-Unlike DPR-based methods that require initial grounding on human written references to retrieve relevant passages, AutoRev employs a GNN that learns to identify important content directly from the paper’s graphical representation, without the need for grounding on existing reviews.
-This not only streamlines the process but also reduces computational overhead while enhancing review quality.
-Our experiments on the ICLR 2024 dataset demonstrate that AutoRev consistently outperforms existing baselines, by 58.72% on popular evaluation metrics like ROUGE, and BERTScore.
-These findings are further strongly backed by human evaluations, where the generated reviews were rated nearly as constructive and acceptable as human-written ones, suggesting AutoRev could meaningfully reduce reviewer workload.
-Beyond these performance gains, we emphasize AutoRev’s broader role in the web ecosystem.
-Academic research papers represent some of the most authoritative and trusted content on the web.
-By accelerating the peer-review workflow and enabling scalable, automated feedback, AutoRev contributes to the timely dissemination of credible scholarly knowledge online.
-In doing so, it supports not only authors and reviewers but also the wider web community that depends on accurate, trustworthy research for discovery, indexing, and downstream applications.
-Future work will explore extending this framework to a broader range of academic domains and investigating further improvements in graph construction and neural retrieval techniques to continue advancing the state-of-the-art in automated review generation.
-Appendix A Ablation studies
-The quantitative performance of different combinations of $k$ (number of review sentences concatenated; $k=1,3,5$) and $m$ (number of top DPR-retrieved passages used to train the GNN; $m=3,5$) for review generation is reported in Table 10 and Table 11, evaluated against the gold-standard reviews.
+Appendix B Ablation studies
+The quantitative performance of different combinations of $k$ (number of review sentences concatenated; $k=1,3,5$) and $m$ (number of top DPR-retrieved passages used to train the GNN; $m=3,5$) for review generation is reported in Table 9 and Table 10, evaluated against the reference feedback.
 While calculating the ROUGE metrics, we enable the use_stemmer option to ensure consistent matching of words with different morphological forms, to improve the reliability of the evaluation.
-Table 10.
-Overall performance of Mistral across different $(k,m)$ settings.
-Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours).
-Bold highlights the best scores, and underlined values indicate the second-best.
 Model ROUGE-F1 ROUGE-Recall BERTScore F1 Training Time (hrs)
@@
 Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
-Table 11.
-Overall performance of Llama across different $(k,m)$ settings.
+Table 9: Overall performance of Mistral across different $(k,m)$ settings.
 Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours).
@@
 Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
-Appendix B GNN Implementation Details
+Table 10: Overall performance of Llama across different $(k,m)$ settings.
+Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours).
+Bold highlights the best scores, and underlined values indicate the second-best.
+Appendix C GNN Implementation Details
+The passing of information from one node to another is enhanced by the GATConv (Equations 2, 3, and 4) layers, which involves the concept of attention to selectively propagate messages in the Message Passing layer of a GNN.
+$\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$ (2)
+$e_{ij}=\text{LeakyReLU}\Bigl(\mathbf{a}^{T}\bigl[\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}\bigr]\Bigr)$ (3)
+$\alpha_{ij}=\frac{\exp\bigl(e_{ij}\bigr)}{\sum_{k\in\mathcal{N}(i)}\exp\bigl(e_{ik}\bigr)}$ (4)
 We train and test our model on a single NVIDIA RTX6000 GPU with 48GB of VRAM.
-B.1. Model Architecture
+C.1 Model Architecture
 The model contains a Graph Neural Network (GNN) module and a passage classifier.
@@
 This architecture enables the model to learn nonlinear transformations of the graph-encoded passage representations while incorporating regularization and normalization for stable training.
-B.2. Training
+C.2 Training
 The model is trained for 30 epochs, with the best validation performance checkpoint retained for inference. The loss function used for training is the Cross Entropy Loss.
-Appendix C Finetuning Implementation Details
+Appendix D Finetuning Implementation Details
 We fine-tune Llama-3.1-8B using two NVIDIA L40S GPUs and Mistral using a single NVIDIA RTX 6000 Ada GPU. Both GPUs are equipped with 48GB of GDDR6 VRAM.
@@
 The full set of hyperparameters used during training is listed in Table 11.
-Table 12. Hyperparameters Used During Finetuning
 Hyperparameter Value
@@
 LR Scheduler Type cosine
-C.1. Prompt for Finetuning
-The prompt used for fine-tuning both Llama 3.1 8B and Mistral-Instruct-v0.2 is shown in Listing C.1.
+Table 11: Hyperparameters Used During Finetuning
+D.1 Prompt for Finetuning
+The prompt used for fine-tuning both Llama 3.1 8B and Mistral-Instruct-v0.2 is shown in Listing D.1.
 […]
-The extracted relevant passages are inserted ahead of the ”Research Paper Passages” heading.
-Similarly, the corresponding content for each review section (Summary, Strengths, Weaknesses, and Questions) is added after their respective headings in the generated review.
-Appendix D Human Evaluation
-Due to space constraints, the detailed evaluation guidelines are made available in an anonymized GitHub repository: Anonymous Repository.
-For the camera-ready version, we will release these materials in a public GitHub repository to ensure full transparency and accessibility.
-We will now proceed to show the snapshots of our annotation UI, developed using Streamlit  [^10] https://streamlit.io/.
-D.1. Qualitative Assessment
-Examples of the user interface used for the Qualitative Assessment can be seen in Figure 4 and Figure 5.
-Figure 4.
-This interface component corresponds to the Qualitative Assessment task described in Section 5.3.1.
-It marks the beginning of the review presentation.
-Due to space constraints, only the summary and a subset of the strengths from one review are shown here.:
-Figure 5. This interface component corresponds to the task described in Section 5.3.1. In this part, annotators provide a final rating across the four categories and submit their annotation.:
-D.2. Factual Validity Assessment
-Examples of the user interface used for Factual Validity Assessment can be seen in Figure 6a and Figure 6b.
+The extracted relevant passages are inserted ahead of the "Research Paper Passages" heading.
+Similarly, the corresponding content for each feedback section (Summary, Strengths, Weaknesses, and Questions) is added after their respective headings in the generated feedback.
+Appendix E LLM-as-a-judge and Qualitative Assessment
+The prompt used for LLM-as-a-judge with Gemini-2.0-Flash is shown in Figure 4. The corresponding model-generated feedback is inserted under the appropriate subheadings within the prompt.
+For Qualitative Assessment, the prompt in Figure 4 is modified only slightly as to just remove the output format.
+This slightly modified prompt is given as the evaluation guideline to the annotators.
+It is ensured that the annotators selected are eligible to review papers for the particular conference.
+** Instruction ** : You are an expert on research papers. Evaluate the following feedback according to the scoring
+[…]
+Give rating for the categories described below on the full scale from 1(lowest rating) to 5(highest rating).
+Assign scores independently per category (they may differ significantly).
+**Category 1: Feedback Depth**
+1 – Unqualified: Admits lack of expertise | Fundamental misunderstandings | Fails to engage with technical content
+2 – Limited: Surface-level comprehension | Misses >2 key concepts | Confuses established methods
+3 – Functional: Grasps main contributions | Identifies 1-2 technical flaws | Limited literature context
+4 – Qualified: Precise technical critique | Contextualizes in recent work (>3 citations) |
+Identifies methodological tradeoffs
+5 – Authoritative: Demonstrates field leadership | Predicts follow-up research directions |
+Reveals fundamental limitations in approach
+**Category 2: Feedback Thoroughness**
+1 – Negligent: <3 substantive comments | Skips methodology/results | Fails to address claims
+2 – Patchy: Only addresses obvious aspects | 0 analysis of experiments | No ethical consideration
+3 – Standard: Checks main claims | Basic methodology feedback | 1 reproducibility concern noted
+4 – Rigorous: Line-by-line methodology check | Statistical validity analysis | 2+ improvement suggestions
+5 – Forensic: Alternative approach proposals | Re-analysis of key results |
+Compliance checklist (ethics, reproducibility, etc.)
+**Category 3: Feedback Constructiveness**
+1 – Destructive: Personal attacks | Unsubstantiated dismissal | 0 actionable feedback
+2 – Demanding: Unrealistic requests ("redo all experiments") | Vague quality complaints
+3 – Directive: General improvement areas | 1-2 specific examples | Mixed positive/negative tone
+4 – Strategic: Priority-ranked suggestions | Alternative methodology paths | Conference-specific improvements
+5 – Transformative: Step-by-step revision roadmap | Code/experiment snippets provided | Before-after examples
+**Category 4: Feedback Helpfulness**
+1 – Misleading: Factually incorrect suggestions | Contradicts paper’s goals | Harmful recommendations
+2 – Abstract: "Improve writing" without examples | "Add experiments" without design
+3 – Guided: Section-specific edits | 2-3 implementable suggestions | Partial error diagnosis
+4 – Impactful: Directly address weaknesses | Templates/examples provided | Clear priority levels
+5 – Revelatory: Exposes fundamental flaws | Provides benchmark comparisons | Supplies missing literature
+**Scoring Methodology**
+For EACH category:
+1. Find the HIGHEST level (5-1) where ALL criteria are FULLY satisfied (Base Level)
+2. Check criteria in the NEXT HIGHER LEVEL (Base Level + 1):
+   - Calculate partial credit: (Satisfied Criteria in Higher Level / Total Criteria in Higher Level)
+3. Final Score = Base Level + Partial Credit
+4. Round to 2 decimal places (0.33 increments for 3-criteria levels, 0.5 for 2-criteria levels)
+[…]
+[…]
+Respond ONLY with this JSON format:
+  "feedback_scores": {{
+    "Category 1": "",
+    "Category 2": "",
+    "Category 3": "",
+    "Category 4": ""
+Figure 4: Prompt used during evaluating the quality of generated feedbacks with LLM as a Judge
+Appendix F Human Evaluation UI
+Our annotation UI is developed using Streamlit  [^4] https://streamlit.io/.
+F.1 Qualitative Assessment
+Examples of the user interface used for the Qualitative Assessment are shown in Figure 5 and  6.
+Figure 5: This interface component corresponds to the Qualitative Assessment task described in Section 6.3.1.
+It marks the beginning of the feedback presentation.
+Due to space constraints, only the summary and a subset of the strengths from one feedback are shown here.
+Figure 6: This interface component corresponds to the task described in Section 6.3.1. In this part, annotators provide a final rating across the four categories and submit their annotation.
+F.2 Factual Validity Assessment
+Dataset Feedback Positive Rating (%)
+ICLR 2024 ICT_10 76.97
+Llama $(5,3)$ 65.25
+Llama $(5,5)$ 64.58
+Other Llama $(5,3)$ 61.86
+Llama $(5,5)$ 64.31
+Table 12: Overall Positive Rating Percentage (sum of “Completely Agree” and “Mostly Agree”) for test dataset of ICLR 2024 as well as NeurIPS and COLM 2025 (Other).
+Examples of the user interface used for Factual Validity Assessment can be seen in Figures 7a and  7b.
 (a) The original paper is displayed in a PDF viewer to provide factual context.
 (b) Each point of the review is rated for accuracy using a dropdown menu.
-Figure 6.
-The user interface for the Factual Validity Assessment task described in Section 5.3.2.
+Figure 7: The user interface for the Factual Validity Assessment task described in Section 6.3.2.
 On the left (a), the full research paper is presented to the annotator.
-On the right (b), the annotator rates each point of the review based on its consistency with the source document.:
-Appendix E Example of a Generated Review
-We present a review generated by our system for the paper: On the Role of Edge Dependency in Graph Generative Models [^11] https://openreview.net/forum?id=LCQ7YTzgRQ
+On the right (b), the annotator rates each point of the feedback based on its consistency with the source document.
+Appendix G Evaluation Guidelines for Factual Validity Assessment
+The evaluation guidelines for the factual validity assessment are shown in Figure 8. It is ensured that the annotators selected are eligible to review papers for the particular conference.
+**Instructions**:
+To assess a structured feedback and check how accurately it reflects the paper’s content.
+**Evaluation Process**:
+1. Read the entire paper first.
+   - If you need to download the paper for some reason, do ensure that you only download the paper using the provided UI.
+2. Read the structured feedback and identify every DISCRETE CLAIM.
+   - For the “Summary” section of the review, treat the entire content of the Summary
+     as ONE discrete claim.
+   - For the “Strengths”, “Weaknesses”, and “Questions” sections of the review,
+     treat EACH bulleted point as a separate discrete claim.
+3. For each claim, choose the best category from the dropdown.
+**Claim Categorization**:
+You must rate the Validity Agreement of every claim on a scale of 1–4,
+where 1 indicates “Completely Disagree” and 4 indicates “Completely Agree”.
+**Completely Disagree (Invalid Claim) [1]**:
+- The claim is factually wrong, contradicts the paper, or is hallucinated.
+**Mostly Disagree (Largely Invalid Claim) [2]**:
+- The claim is mostly incorrect, though it may contain minor truth.
+**Mostly Agree (Largely Valid Claim) [3]**:
+- The claim is mostly correct, with minor gaps or oversights.
+**Completely Agree (Fully Valid Claim) [4]**:
+- The claim is entirely accurate.
+Figure 8: Guidelines used for human evaluation of factual validity of the structured feedback
+Appendix H Example of a Generated Structured Feedback
+We present a structured feedback generated by our system in Figure 5
+On the Role of Edge Dependency in Graph Generative Models
 Summary
-The paper introduces a novel framework for evaluating generative models of graphs, focusing on the concept of “overlap” to ensure both accuracy and edge diversity.
-It categorizes graph generative models into three levels of complexity: edge independent, node independent, and fully dependent models.
-The paper derives theoretical bounds on triangle counts and other short-length cycles based on overlap and provides new generative models for each level.
-However, the paper’s novelty and practical contributions are questioned due to the use of existing methods and the lack of significant empirical validation.
-The experimental section is seen as limited and does not convincingly demonstrate the superiority of the proposed methods over existing ones.
+The paper introduces a novel framework for evaluating generative models of graphs,
+focusing on the concept of ‘‘overlap’’ to ensure both accuracy and edge diversity.
+It categorizes graph generative models into three levels of complexity: edge
+independent, node independent, and fully dependent models. The paper derives
+theoretical bounds on triangle counts and other short-length cycles based on overlap
+and provides new generative models for each level. However, the paper’s novelty and
+practical contributions are questioned due to the use of existing methods and the
+lack of significant empirical validation. The experimental section is seen as
+limited and does not convincingly demonstrate the superiority of the proposed
+methods over existing ones.
 Strengths
-  • The paper introduces a novel evaluation framework that considers the importance of model-generated graph overlap to ensure both accuracy and edge diversity, which is a significant contribution to the field of graph generative models.
-  • The paper presents a hierarchy of graph generative models categorized into three levels of complexity, which encapsulates a wide range of prevalent methods and provides a structured framework for evaluating graph generative models.
-  • The theoretical bounds on the number of triangles and other short-length cycles are well-presented and support the findings with instances demonstrating the asymptotic optimality of these bounds.
-  • The authors introduce new generative models for each of the three hierarchical levels, leveraging dense subgraph discovery (Gionis & Tsourakakis, 2015), which is a practical application of the theoretical findings.
+- The paper introduces a novel evaluation framework that considers the importance of
+  model-generated graph overlap to ensure both accuracy and edge diversity, which is
+  a significant contribution to the field of graph generative models.
+- The paper presents a hierarchy of graph generative models categorized into three
+  levels of complexity, which encapsulates a wide range of prevalent methods and
+  provides a structured framework for evaluating graph generative models.
+- The theoretical bounds on the number of triangles and other short-length cycles are
+  well-presented and support the findings with instances demonstrating the asymptotic
+  optimality of these bounds.
+- The authors introduce new generative models for each of the three hierarchical
+  levels, leveraging dense subgraph discovery (Gionis & Tsourakakis, 2015), which
+  is a practical application of the theoretical findings.
 Weaknesses
-  • The novelty of the paper is limited as it primarily relies on existing methods and theoretical frameworks, and the experimental results do not convincingly demonstrate the superiority of the proposed methods over existing ones.
-  • The empirical evaluation is weak and does not provide sufficient evidence to support the claims made in the paper.
-The experimental results are not robust and the paper does not provide enough comparisons with other models to establish the effectiveness of the proposed methods.
-  • The paper has several presentation issues, including unclear definitions and notations, which could hinder the understanding of the content for readers not familiar with graph generative models.
-  • The practical applications and the significance of the proposed methods are not clearly demonstrated, and the paper lacks a detailed discussion on the limitations and potential future work.
+- The novelty of the paper is limited as it primarily relies on existing methods and
+  theoretical frameworks, and the experimental results do not convincingly
+  demonstrate the superiority of the proposed methods over existing ones.
+- The empirical evaluation is weak and does not provide sufficient evidence to
+  support the claims made in the paper. The experimental results are not robust and
+  the paper does not provide enough comparisons with other models to establish the
+  effectiveness of the proposed methods.
+- The paper has several presentation issues, including unclear definitions and
+  notations, which could hinder the understanding of the content for readers not
+  familiar with graph generative models.
+- The practical applications and the significance of the proposed methods are not
+  clearly demonstrated, and the paper lacks a detailed discussion on the limitations
+  and potential future work.
 Questions
-  • Can the authors clarify the definitions and notations used in the paper, particularly the ones related to the overlap and triangle density, to ensure a clearer understanding of the content?
-  • How do the authors justify the novelty of their work given the reliance on existing methods and theoretical frameworks?
-  • What specific contributions does the paper make to the field of graph generative models, and can these be clearly articulated?
-  • Why was the empirical evaluation conducted in a limited manner, and how can the paper be improved to provide more robust and convincing experimental results?
-  • Could the authors discuss the practical applications of the proposed methods and the significance of their work in more detail?
-  • Are there any plans to extend the theoretical bounds to higher-order cycles, and how might this affect the overall conclusions drawn in the paper?
+- Can the authors clarify the definitions and notations used in the paper,
+  particularly the ones related to the overlap and triangle density, to ensure a
+  clearer understanding of the content?
+- How do the authors justify the novelty of their work given the reliance on
+  existing methods and theoretical frameworks?
+- What specific contributions does the paper make to the field of graph generative
+  models, and can these be clearly articulated?
+- Why was the empirical evaluation conducted in a limited manner, and how can the
+  paper be improved to provide more robust and convincing experimental results?
+- Could the authors discuss the practical applications of the proposed methods and
+  the significance of their work in more detail?
+- Are there any plans to extend the theoretical bounds to higher-order cycles, and
+  how might this affect the overall conclusions drawn in the paper?
+Figure 9: Output of the generated feedback by AutoRev for the paper On the Role of Edge Dependency in Graph Generative Models [^5] https://openreview.net/forum?id=LCQ7YTzgRQ